### v1

# Communication-Efficient Stochastic Distributed Learning

Abstract
We address distributed learning problems, both nonconvex and convex, over undirected networks. In particular, we design a novel algorithm based on the distributed Alternating Direction Method of Multipliers (ADMM) to address the challenges of high communication costs, and large datasets. Our design tackles these challenges i) by enabling the agents to perform multiple local training steps between each round of communications; and ii) by allowing the agents to employ stochastic gradients while carrying out local computations. We show that the proposed algorithm converges to a neighborhood of a stationary point, for nonconvex problems, and of an optimal point, for convex problems. We also propose a variant of the algorithm to incorporate variance reduction thus achieving exact convergence. We show that the resulting algorithm indeed converges to a stationary (or optimal) point, and moreover that local training accelerates convergence. We thoroughly compare the proposed algorithms with the state of the art, both theoretically and through numerical results.
Index Terms:
Distributed learning; Stochastic optimization; Variance reduction; Local training.
I Introduction
Recent technological advances have enabled the widespread adoption of devices with computational and communication capabilities in many fields, for instance, power grids [1], robotics [2, 3], transportation networks [4], and sensor networks [5]. These devices connect with each other, forming multi-agent systems that cooperate to collect and process data [6]. As a result, there is a growing need for algorithms that enable efficient and accurate cooperative learning.
In specific terms, the objective in distributed learning is to train a model (e.g., a neural network) with parameters $x\in\mathbb{R}^{n}$ cooperatively across a network of $N$ agents. Each agent has access to a local dataset which defines the local cost as
$f_{i}({x})=\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}f_{i,h}(x)\,,$ (1)
with $f_{i,h}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ being the loss function associated to data point $h\in\{1,\ldots,m_{i}\}$. Thus, the goal is for the agents to solve the following constrained problem [7, 8]:
$\min_{x_{i}\in\mathbb{R}^{n},\ i\in\mathcal{V}}\ \ \frac{1}{N}\sum_{i=1}^{N}f_{i}({x}_{i})\quad\text{s.t.}\ \ x_{1}=x_{2}=\cdots=x_{N}\,,$ (2)
where the objective is the sum of local costs (1) to pool together the agents’ data. Moreover, each agent is assigned a set $x_{i}$ of local model parameters, and the consensus constraints $x_{1}=x_{2}=\ldots=x_{N}$ ensure that the agents will asymptotically agree on a shared trained model.
TABLE I: Comparison with the state of the art in stochastic distributed optimization.
Algorithm [Ref.] variance reduction grad. steps $\div$ comm. # stored variables comm. size # $\nabla f_{i,j}$ evaluations per iteration assumpt. convergence
K-GT [9] $\tau\div 1$ $2$ $2|\mathcal{N}_{i}|$ $1$ n.c. sub-linear, $\propto\sigma^{2}$
LED [10] $\tau\div 1$ $2$ $|\mathcal{N}_{i}|$ $1$ n.c. s.c. sub-linear, $\propto\sigma^{2}$ linear, $\propto\sigma^{2}$
RandCom [11] $\left\lceil\frac{1}{p}\right\rceil\div 1$ (in mean) $2$ $|\mathcal{N}_{i}|$ $1$ n.c. s.c. sub-linear, $\propto\sigma^{2}$ linear, $\propto\sigma^{2}$
VR-EXTRA/DIGing [12], GT-VR [13] $1\div 1$ $3$ $2|\mathcal{N}_{i}|$ $|\mathcal{B}|$, $m_{i}$ every $\left\lceil\frac{1}{p}\right\rceil$ n.c. s.c. sub-linear, $\to 0$ linear, $\to 0$
GT-SAGA [14, 15] $1\div 1$ $3$ $2|\mathcal{N}_{i}|$ $1$ n.c. s.c. sub-linear, $\to 0$ linear, $\to 0$
GT-SARAH [16] $1\div 1$ $3$ $2|\mathcal{N}_{i}|$ $|\mathcal{B}|$, $m_{i}$ every $\tau$ n.c. sub-linear, $\to 0$
GT-SVRG [14] $1\div 1$ $3$ $2|\mathcal{N}_{i}|$ $1$, $m_{i}$ every $\tau$ s.c. linear, $\to 0$
LT-ADMM [this work] $\tau\div 1$ $|\mathcal{N}_{i}|+1$ $|\mathcal{N}_{i}|$ $|\mathcal{B}|$ n.c. sub-linear, $\propto\sigma^{2}$
LT-ADMM-VR [this work] $\tau\div 1$ $|\mathcal{N}_{i}|+1$ $|\mathcal{N}_{i}|$ $|\mathcal{B}|$, $m_{i}$ every $\tau$ n.c. sub-linear, $\to 0$
number of vectors in $\mathbb{R}^{n}$ stored by each agent between iterations (disregarding temporary variables)
number of messages sent by each agent during a communication round
n.c. and s.c. stand for (non)convex and strongly convex
To effectively tackle this problem, especially when dealing with large datasets that involve sensitive information, distributed methods have become increasingly important. These techniques offer significant robustness advantages over federated learning algorithms [6], as they do not rely on a central coordinator and thus, for example, have not a single point of failure. In particular, both distributed gradient-based algorithms [17, 18, 19, 20], and distributed Alternating Direction Method of Multipliers (ADMM) [21, 22, 23, 24] have proven to be effective strategies for solving such problems.
However, many learning applications face the challenges of: high communication costs, especially when training large models, and large datasets. In this paper, we jointly address these challenges with the following approach. First, we guarantee the communication efficiency of our algorithm by adopting the paradigm of local training, which reduces the frequency of communications. In other terms, the agents perform multiple training steps between communication rounds. We tackle the second challenge by locally incorporating stochastic gradients. The idea is to allow the agents to estimate local gradients by employing only a (random) subset of the available data, thus avoiding the computational burden of full gradient evaluations on large datasets.
Our main contributions are as follows:
  • We propose two algorithms based on distributed ADMM, with one round of communication between multiple local update steps. The first algorithm, Local Training ADMM (LT-ADMM), uses stochastic gradient descent (SGD) for the local updates, while the second algorithm, LT-ADMM with Variance Reduction (LT-ADMM-VR), uses a variance-reduced SGD method [25].
  • We establish the convergence properties of LT-ADMM for both nonconvex and convex (not strongly convex) learning problems. In particular, we show almost-sure and mean-squared convergence of LT-ADMM to a neighborhood of the stationary point in the nonconvex case, and to a neighborhood of an optimum in the convex case. The radius of the neighborhood depends on specific properties of the problem and on tunable parameters. We prove that the algorithm achieves a convergence rate of $\mathcal{O}(\frac{1}{K\tau})$, where $K$ is the number of iterations, and $\tau$ the number of local training steps.
  • For LT-ADMM-VR, we prove exact convergence to a stationary point in the nonconvex case, and to an optimum in the convex case. The algorithm has a $\mathcal{O}(\frac{1}{K\tau})$ rate of convergence, which is faster than $\mathcal{O}(\frac{1}{K})$ obtained by related algorithms [16, 15, 13].
  • We provide extensive numerical evaluations comparing the proposed algorithms with the state of the art. The results validate the communication efficiency of the algorithms. Indeed, LT-ADMM and LT-ADMM-VR outperform alternative methods when communications are expensive.
I-A Comparison with the State of the Art
We compare our proposed algorithms – LT-ADMM and LT-ADMM-VR– with the state of the art. The comparison is holistically summarized in Table I.
Decentralized learning algorithms, as first highlighted in the seminal paper [26] on federated learning, face the fundamental challenge of high communication costs. The authors of [26] address this challenge by designing a communication-efficient algorithms which allows the agents to perform multiple local training steps before each round of communication with the coordinator. However, the accuracy of the algorithm in [26] degrades significantly when the agents have heterogeneous data. Since then, alternative federated learning algorithms, e.g., [27, 28, 29, 30], have been designed to employ local training without compromising accuracy. The interest for communication-efficient algorithms has more recently extended to the distributed set-up, where agents rely on peer-to-peer communications rather than on a coordinator, as in federated learning. Distributed algorithms with local training have been proposed in [31, 9, 10, 11]. In particular, [31, 9, 10] present gradient tracking methods which allow each agent to perform a fixed number of local updates between each communication round. The algorithm in [11], which builds on [28], instead triggers communication rounds according to a given probability distribution, resulting in a time-varying number of local training steps. Another related algorithm is that of [32], which allows for both multiple consensus and gradient steps in each iteration. However, this algorithm requires a monotonically increasing number of communication rounds in order to guarantee exact convergence. A stochastic version of [32] was then studied in [33]. The algorithm has inexact gradient evaluations, but only allows for multiple consensus steps.
When the agents employ stochastic gradients in the algorithms of [9, 10, 11], they only converge to a neighborhood of a stationary point, whose radius is proportional to the stochastic gradient variance. Different variance reduction techniques are available to improve the accuracy of (centralized) algorithms relying on stochastic gradients, e.g., [34, 25, 35]. Then, these methods have been applied to distributed optimization by combining them with widely used gradient tracking algorithms [12, 13, 14, 15, 16]. The resulting algorithms succeed in guaranteeing exact convergence to a stationary point despite the presence of gradient noise. However, they are not communication-efficient, as they only allow one gradient update per communication round.
We conclude by providing a summary of the features of the algorithms compared so far in Table I. First, we classify them based on whether they use or not variance reduction and local training. For the latter, we report the ratio of gradient steps to communication rounds that characterizes each algorithm, with a ratio of $1\div 1$ signifying that no local training is used. Notice that LT-ADMM-VR is the only algorithm to use both variance reduction and local training, while the other only use one technique. We then compare the number of variables stored by the agents when they deploy each algorithm (disregarding temporary variables). We notice that the variable storage of LT-ADMM and LT-ADMM-VR, differently from the alternatives, scales with the size of an agent’s neighborhood; this is due to the use of distributed ADMM as the foundation of our proposed algorithms [21]. We see that [10, 11], LT-ADMM, and LT-ADMM-VR require one communication per neighbor, while the other methods require two communications per neighbor. We also compare the algorithms by the computational complexity of the gradient estimators they employ, namely, the number of component gradient evaluations needed per local training iteration. The algorithms of [9, 10, 11, 14] use a single data point to estimate the gradient, while [12, 16], LT-ADMM, LT-ADMM-VR can apply mini-batch estimators that use a subset $\mathcal{B}$ of the local data points. The use of mini-batches yields more precise gradient estimates and increased flexibility. However, we remark that the gradient estimators used in [12, 16, 14], LT-ADMM, LT-ADMM-VR require a registry of component gradient evaluations, which needs to be refreshed entirely at fixed intervals. This coincides with the evaluation of a full gradient, and thus requires $m_{i}$ component gradient evaluations. Finally, we compare the algorithms’ convergence. We notice that all algorithms, except for [14], provide (sub-linear) convergence guarantees for convex and nonconvex problems. Additionally, some works show linear convergence for strongly convex problems. We further distinguish between algorithms which achieve exact convergence due to the use of variance reduction, or inexact convergence with an error proportional to the stochastic gradient variance ($\propto\sigma^{2}$).
Outline
The outline of the paper is as follows. Section II formulates the problem at hand, and presents the proposed algorithms design. Section III analyzes their convergence, and discusses the results. Section IV reports and discusses numerical results comparing the proposed algorithms with the state of the art. Section V presents some concluding remarks.
Notation
[…]
II Problem Formulation and Algorithm Design
In this section, we formulate the problem at hand and present our proposed algorithms.
II-A Problem Formulation
We target the solution of (2) over a (undirected) graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{1,...,N\}$ is the set of $N$ agents, and $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ is the set of edges $(i,j)$, $i,j\in\mathcal{V}$. In particular, we assume that the local costs $f_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ are in the empirical risk minimization form (1). We make the following assumptions for (2), which commonly underpin the convergence analysis of distributed algorithms.
Assumption 1.
$\mathit{\mathcal{G}}=(\mathcal{V},\mathcal{E})$ is a connected, undirected graph.
Assumption 2.
The cost function $f_{i}$ of each agent $i\in\mathcal{V}$ is $L$-smooth. That is, there exists $l>0$ such that $\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|$, $\forall x,y\in\mathbb{R}^{n}$. Moreover, $f_{i}$ is proper: $f_{i}(x)>-\infty$, $\forall x\in\mathbb{R}^{n}$.
When, in the following, we specialize our results to convex scenarios, we resort to the further assumption below.
Assumption 3.
Each function $f_{i}$, $i\in\mathcal{V}$, is convex.
II-B Algorithm Design
We start our design from the distributed ADMM, characterized by the updates [^1] We remark that, more precisely, (3) corresponds to the algorithm in [21] with relaxation parameter $\alpha=1/2$. [21]:
$x_{i,k+1}$ $=\operatorname{prox}_{f_{i}}^{1/\rho\left|\mathcal{N}_{i}\right|}\Bigg{(}\frac{1}{\rho\left|\mathcal{N}_{i}\right|}\sum_{j\in\mathcal{N}_{i}}z_{ij,k}\Bigg{)}\,,$ (3a)
$z_{ij,k+1}$ $=\frac{1}{2}\left(z_{ij,k}-z_{ji,k}-2\rho x_{j,k+1}\right)\,,$ (3b)
where $\mathcal{N}_{i}=\{j\in\mathcal{V}\ |\ (i,j)\in\mathcal{E}\}$ denotes the neighbors of agent $i$, $\rho>0$ is a penalty parameter, and $z_{ij}\in\mathbb{R}^{n}$ are auxiliary variables, one for each neighbor of agent $i$. This algorithm converges in a wide range of scenarios, and, differently from most gradient tracking approaches, shows robustness to many challenges (asynchrony, limited communications, etc) [21, 22]. However, the drawback of (3) is that the agents need to solve an optimization problem to update $x_{i}$, which in general does not have a closed-form solution. Therefore, in practice, the agents need to compute an approximate update (3a), which can lead to inexact convergence [22].
In this paper, we modify (3) to use approximate local updates, while ensuring that this choice does not compromise exact convergence. In particular, we allow the agents to use $\tau\in\mathbb{N}$ iterations of a gradient-based solver to approximate (3a), which yields the update:
$\begin{split}&\phi_{i,k}^{0}=x_{i,k}\,,\\ &\phi_{i,k}^{t+1}=\phi_{i}^{t}-\gamma\Bigg{(}g_{i}(\phi_{i,k}^{t})+\rho\left|\mathcal{N}_{i}\right|\phi_{i,k}^{t}-\sum_{j\in\mathcal{N}_{i}}z_{ij,k}\Bigg{)},\\ &\qquad t=0,\ldots,\tau-1\,,\\ &x_{i,k+1}=\phi_{i,k}^{\tau}\,,\end{split}$ (4)
where $g_{i}(\phi_{i,k}^{\ell})$ is an estimate of the gradient $\nabla f_{i}$. The resulting algorithm is a distributed gradient method, with the difference that each communication round (3b) is preceded by $\tau>1$ local gradient evaluations. This is an application of the local training paradigm [10]. We remark that the convergence of the proposed algorithm rests on the initialization $\phi_{i,k}^{0}=x_{i,k}$, which enacts a feedback loop on the local training. In general, without this initialization, exact convergence cannot be achieved [22].
The local training (4) requires a local gradient evaluation or at least its estimate. In the following, we introduce two different estimator options. Notice that the gradient of the penalty term, $\rho\left|\mathcal{N}_{i}\right|\phi_{i,k}^{t}-\sum_{j\in\mathcal{N}_{i}}z_{ij,k}$, is exactly known and does not need an estimator. The most straightforward idea is to simply employ a local gradient $g_{i}(\phi)=\nabla f_{i}(\phi)$. However, in learning applications, the agents may store large datasets ($m_{i}\gg 1$). Therefore, computing $\nabla f_{i}(\phi)$ becomes computationally expensive. To remedy this, the agents can instead use stochastic gradients, choosing
$g_{i}(\phi)=\frac{1}{|\mathcal{B}_{i}|}\sum_{h\in\mathcal{B}_{i}}\nabla f_{i,h}(\phi)\,,$ (5)
where $\mathcal{B}_{i}$ are randomly drawn indices from $\{1,\ldots,m_{i}\}$, with $|\mathcal{B}_{i}|<m_{i}$. While reducing the computational complexity of the local training iterations, the use of stochastic gradients results in inexact convergence. The idea therefore is to employ a gradient estimator based on a variance reduction scheme. In particular, we use the scheme proposed in [25], which demonstrates faster convergence in practice and is characterized by the following procedure. Each agent maintains a table of component gradients $\{\nabla f_{i,h}(r_{i,h,k}^{t})\},h=1,\ldots,m_{i}$, where $r_{i,h,k}^{t}$ is the most recent iterate at which the component gradient was evaluated. This table is reset at the beginning of every new local training (that is, for any $k\in\mathbb{N}$ when $t=0$). Using the table, the agents then estimate their local gradients as
$\begin{split}g_{i}\left(\phi_{i,k}^{t}\right)&=\frac{1}{|\mathcal{B}_{i}|}\sum_{h\in\mathcal{B}_{i}}\left(\nabla f_{i,h}\left(\phi_{i,k}^{t}\right)-\nabla f_{i,h}\left(r_{i,h,k}^{t}\right)\right)\\ &+\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\nabla f_{i,h}(r_{i,h,k}^{t}).\end{split}$ (6)
The gradient estimate is then used to update $\phi_{i,k}^{t+1}$ according to (4); afterwards, the agents update their local memory by setting $r_{i,h,k}^{t+1}=\phi_{i,k}^{t+1}$ if $h\in\mathcal{B}_{i}$, and $r_{i,h,k}^{t+1}=r_{i,h,k}^{t}$ otherwise. Notice that this update requires a full gradient evaluation at the beginning of each local training, to populate the memory with $\{\nabla f_{i,h}(r_{i,h,k}^{0})=\nabla f_{i,h}(\phi_{i,k}^{0})\},h=1,\ldots,m_{i}$. In the following steps ($t>0$), each agent only computes $|\mathcal{B}_{i}|$ component gradients.
The resulting algorithms are reported in Algorithm 1.
Algorithm 1 LT-ADMM and LT-ADMM-VR
1:For each node $i$, initialize $x_{i,0}=z_{ij,0}$, $j\in\mathcal{N}_{i}$. Set the penalty $\rho$, the number of local training steps $\tau$, and the local step-size $\gamma$.
2:for $k=0,1,\ldots$ every agent $i$ do
3:     // local training
4:     $\phi_{i,k}^{0}=x_{i,k}$, $r_{i,h,k}^{0}=x_{i,k}$, for all $h\in\{1,\ldots.m_{i}\}$
5:     for $t=0,1,\ldots,\tau-1$ do
6:         draw the batch $\mathcal{B}_{i}$ uniformly at random
7:         update the gradient estimator according to (5)
8:         update the gradient estimator according to (6)
9:         update $\phi_{i,k}$ according to (4)
10:         if $h\in\mathcal{B}_{i}$ update $r_{i,h,k}^{t+1}=\phi_{i,k}^{t+1}$, else $r_{i,h,k}^{t+1}=r_{i,h,k}^{t}$
11:     end for
12:     $x_{i,k+1}=\phi_{i,k}^{\tau}$
13:     // communication
14:     transmit $z_{ij,k}-2\rho x_{i,k+1}$ to each neighbor $j\in\mathcal{N}_{i}$, and receive the corresponding transmissions
15:     // auxiliary update
16:     update $z_{ij,k+1}$ according to (3b)
17:end for
III Convergence Analysis and Discussion
In this section, we analyze the convergence rate of Algorithm 1 in both nonconvex and convex scenarios. Throughout, we will employ the following metric of convergence
$\mathcal{D}_{k}=\mathbb{E}\left[\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}+\frac{1}{\tau}\sum_{t=0}^{\tau-1}\left\lVert\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\right\rVert^{2}\right],$ (7)
where $F(x)=\frac{1}{N}\sum_{i=1}^{N}f_{i}(x)$ and $\bar{x}_{k}=\frac{1}{N}\sum_{i=1}^{N}x_{i}$. If the agents’ converge to a stationary point of (2), then $\mathcal{D}_{k}\to 0$.
III-A Convergence with SGD
We start by characterizing the convergence of Algorithm 1 when the agents use SGD during local training. To this end, we make the following standard assumption on the variance of the gradient estimators.
Assumption 4.
For all $\phi\in\mathbb{R}^{n}$ the gradient estimators $g_{i}(\phi)$, $i\in\mathcal{V}$, in (5) are unbiased and their variance is bounded by some $\sigma^{2}>0$:
$\begin{gathered}\mathbb{E}\left[g_{i}(\phi)-\nabla f_{i}(\phi)\right]=0\\ \mathbb{E}\left[\left\|g_{i}(\phi)-\nabla f_{i}(\phi)\right\|^{2}\right]\leq\sigma^{2}\end{gathered}.$
We are now ready to state our convergence results. All the proofs are deferred to the Appendix.
Theorem 1 (Nonconvex case).
Let Assumptions 1, 2, and 4 hold. If the local step-size $\gamma$ is chosen sufficiently small (see (10) for the precise bound), then the output of LT-ADMM satisfies:
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\mathcal{O}\left(\frac{F(\bar{x}_{0})-F(x^{*})}{K\gamma\tau}\right)+\mathcal{O}\left(\gamma\tau\sigma^{2}\right)+\mathcal{O}\left(\frac{\|\widehat{\mathbf{d}}_{0}\|^{2}}{\rho KN}\right),$ (8)
where $x^{*}$ is a stationary point of (2), and $\|\widehat{\mathbf{d}}_{0}\|$ is related to the initial conditions (see (26)).
Theorem 1 shows that LT-ADMM converges to a neighborhood of a stationary point $x^{*}$ as $K\to\infty$. The radius of this neighborhood is proportional to the step-size $\gamma$, to the number of local training epochs $\tau$, and to the stochastic gradient variance $\sigma^{2}$. The result can then be particularized to the convex case as follows.
Corollary 1 (Convex case).
In the setting of Theorem 1, with the additional Assumption 3, then the output of LT-ADMM converges to a neighborhood of an optimal solution $x^{*}$ characterized by (8).
Remark 1 (Exact convergence).
Clearly, if we employ full gradients (and thus $\sigma=0$), then these results prove exact convergence to a stationary/optimal point. This verifies that our algorithm design achieves convergence despite the use of approximate local updates.
III-B Convergence with Variance Reduction
The results of the previous sections show that only inexact convergence can be achieved when employing SGD. The following results show how Algorithm 1 achieves exact convergence when using variance reduction.
Theorem 2 (Nonconvex case).
Let Assumptions 1, 2, and 4 hold. If the local step-size $\gamma$ is chosen sufficiently small (see (11) for the precise bound), then the output of LT-ADMM converges to a stationary point $x^{*}$ of (2), and in particular it holds:
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\mathcal{O}\left(\frac{F\left(\bar{x}_{0}\right)-F(x^{*})}{K\gamma\tau}\right)+\mathcal{O}\left(\frac{\|\widehat{\mathbf{d}}_{0}\|^{2}}{\rho^{2}KN}\right)$ (9)
Corollary 2 (Convex case).
In the setting of Theorem 2, with the additional Assumption 3, then the output of LT-ADMM-VR converges to an optimal solution $x^{*}$, with rate characterized by (9).
III-C Discussion
III-C1 Choice of step-size
The upper bounds to the step-sizes of LT-ADMM and LT-ADMM-VR ((10) and (11) in Appendix A-A), highlight a dependence on several features of the problem. In particular, the step-size bounds decrease as the smoothness constant $L$ increases, as is usually the case for gradient-based algorithms. Moreover, the bounds are proportional to the network connectivity, represented by the second-largest eigenvalue of $\mathcal{G}$’s Laplacian ($|\tilde{\lambda}_{\max}|$). Thus, less connected graphs (smaller $|\tilde{\lambda}_{\max}|$) result in smaller bounds. Finally, we remark that the step-size bound for LT-ADMM-VR is proportional to $\frac{m_{l}}{m_{u}}=\frac{\min_{i\in\mathcal{V}}m_{i}}{\max_{i\in\mathcal{V}}m_{i}}$, where $m_{i}$ is the number of data points available to agent $i$ (see (1)). This ratio can be viewed as a measure of heterogeneity between the agents. Smaller values of $\frac{m_{l}}{m_{u}}$ highlight higher imbalance in the amount of data available to the agents. The step-size bound thus is smaller for less balanced scenarios.
The step-size bounds also depend on the tunable parameters $\tau$, the number of local updates, and $\rho$, the penalty parameter. Therefore, these two parameters can be tuned in order to increase the step-size bounds, which translates in faster convergence.
III-C2 Convergence rates comparison
As discussed in section I-A, various distributed algorithms with variance reduction have been recently proposed, for example, [12, 14] for strongly convex problems, and [16, 15, 13] for nonconvex problems. Focusing on [16, 15, 13], we notice that their convergence rate is $\mathcal{O}(\frac{1}{K})$, while Theorem 2 shows that LT-ADMM-VR has rate of $\mathcal{O}(\frac{1}{\tau K})$. This shows that employing local training accelerates convergence.
Similarly to LT-ADMM-VR, [16, 12] also use batch gradient computations, i.e., they update a subset of components to estimate the gradient (see (6)). Interestingly, the step-size upper bound and, hence, the convergence rate in [16, 12] depend on the batch size. On the other hand, our theoretical results are not affected by the batch size, since we use a different variance reduction technique.
III-C3 Choice of variance reduction mechanism
In variance reduction, we distinguish two main classes of algorithms: those that need to periodically (or randomly) perform a full gradient evaluation (SARAH-type [35]), and those that do not (SAGA-type [25]). In distributed learning, SARAH-type algorithms were proposed in e.g., [16, 14], while SAGA-type algorithms in e.g. [14]. Also the proposed LT-ADMM-VR requires a periodic full gradient evaluation, as the agents re-initialize their local gradient memory at the start of each local training (since they set $r_{i,h,k}^{0}=x_{i,k}$). Clearly, periodically computing a full gradient significantly increases the computational complexity of the algorithm. Thus, one can design a SAGA-type variant of LT-ADMM-VR by removing the gradient memory re-initialization at the start of local training (choosing now $r_{i,h,k}^{0}=r_{i,h,k-1}^{\tau}$). This variant is computationally cheaper and shows promising empirical performance, see the results for LT-ADMM-VR v2 in section IV. However, using the outdated gradient memory leads to a more complex theoretical analysis, which we leave for future work.
III-C4 Uncoordinated parameters
In principle, the agents could use uncoordinated parameters, based on the resources available to them. For example, different agents could employ different local solvers and batch sizes. Alternatively, they could use the same solver but with different step-sizes tailored to the smoothness of the local costs.
IV Numerical Results
In this section we compare the proposed algorithms with the state of the art, applying them to a classification problem with nonconvex regularization, characterized by [10]:
$f_{i}(x)=\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\log\left(1+\exp\left(-b_{i,h}a_{i,h}^{\top}x\right)\right)+\epsilon\sum_{\ell=1}^{n}\frac{[x]_{\ell}^{2}}{1+[x]_{\ell}^{2}}$
where $[x]_{\ell}$ is the $\ell$-th component of $x\in\mathbb{R}^{n}$, and $a_{i,h}\in\mathbb{R}^{n}$ and $b_{i,h}\in\{-1,1\}$ are the pairs of feature vector and label, randomly generated with scikit-learn. We choose a ring graph with $N=10$, and set $n=5$, $m_{i}=100$, $\epsilon=0.01$; the initial conditions are randomly chosen as $x_{i,0}\sim\mathcal{N}(0,100I_{n})$. We use stochastic gradients with a batch of $|\mathcal{B}|=1$. All results are averaged over $100$ Monte Carlo iterations. For the algorithms with local training we select $\tau=5$. We also tune the step-sizes of all algorithms to ensure best performance. Finally, as performance metric we employ $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}$, which is zero if the agents have reached a solution of (2).
We compare LT-ADMM and LT-ADMM-VR with local training algorithms LED [10] and K-GT [9], as well as variance reduction algorithms GT-SARAH [16], and GT-SAGA [14]. We also compare with the alternative version LT-ADMM-VR v2 discussed in section III-C3. We account for the computation time (in unspecified units) of each algorithm when scaling the x-axis. This choice deviates from the common approach of using the iteration count for the x-axis, since we are interested in simulating the real performance of these algorithms. In particular, letting $t_{\mathrm{G}}$ be the time for a component gradient evaluation ($\nabla f_{i,h}$), and $t_{\mathrm{C}}$ the time for a round of communications, Table II reports the computation time incurred by each algorithm over the course of $\tau$ iterations.
TABLE II: Computation time of the algorithms over $\tau$ iterations.
Algorithm [Ref.] Time
LED [10] & K-GT [9] $\tau t_{\mathrm{G}}+2t_{\mathrm{C}}$
GT-SARAH [16] $(m_{i}+\tau-1)t_{\mathrm{G}}+2\tau t_{\mathrm{C}}$
GT-SAGA [14] $\tau\left(t_{\mathrm{G}}+2t_{\mathrm{C}}\right)$
LT-ADMM & LT-ADMM-VR v2 $\tau t_{\mathrm{G}}+t_{\mathrm{C}}$
LT-ADMM-VR $(m_{i}+\tau-1)t_{\mathrm{G}}+t_{\mathrm{C}}$
Figure 1 depicts the comparison of the algorithms in Table II in terms of $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}$. The x-axis for each algorithm is scaled to account for the individual computation time, and we plot the results for three choice of the $t_{\mathrm{G}}/t_{\mathrm{C}}$ ratio.
Figure 1: Algorithms’ comparison for different $t_{\mathrm{G}}/t_{\mathrm{C}}$ ratios. The x-axis is in unspecified computation time units to account for the algorithms’ computational complexity.
First of all, we notice that the algorithms without variance reduction (LED, K-GT, LT-ADMM) as expected only converge to a neighborhood of the optimal solution, as stochastic gradients are employed. The asymptotic errors attained by the algorithms are: LED $8.169\times 10^{-4}$, K-GT $7.552\times 10^{-4}$, LT-ADMM $6.327\times 10^{-4}$.
Turning to the algorithms with variance reduction, we see that all of them converge to the optimal solution; however, depending on the value of $t_{\mathrm{G}}/t_{\mathrm{C}}$, their relative speed of convergence changes. When gradient computations are cheaper than communications ($t_{\mathrm{G}}/t_{\mathrm{C}}=0.1$), the proposed algorithm LT-ADMM-VR (and LT-ADMM-VR v2) outperform both GT-SARAH and GT-SAGA, since the latter do not employ local training. This testifies to the benefit of employing local training in scenarios where communications are expensive. As the ratio $t_{\mathrm{G}}/t_{\mathrm{C}}$ increases to $1$ and then $10$, we see how LT-ADMM-VR and GT-SARAH, on the one hand, and LT-ADMM-VR v2 and GT-SAGA, on the other hand, tend to align in terms of performance, as the bulk of the computation time is now due to gradient evaluations, of which the two pairs of algorithms have a similar number (see Table II).
We conclude evaluating the convergence speed of LT-ADMM-VR for different choices of the number of local training epochs $\tau$. Figure 2 reports the computation time (scaled according to Table II) to reach $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}<10^{-9}$.
Figure 2: Computation time needed by LT-ADMM-VR to reach $\left\lVert\nabla F(\bar{x}_{k})\right\rVert^{2}<10^{-9}$ for different numbers of local training epochs $\tau$.
Interestingly, it appears that there is a finite optimal value ($\tau=8$), while smaller and larger values lead to slower convergence.
V Concluding Remarks
In this paper, we considered (non)convex distributed learning problems. In particular, to address the challenge of expensive communication, we proposed two communication-efficient algorithms, LT-ADMM and LT-ADMM-VR, that use local training. The algorithms employ SGD and SGD with variance reduction, respectively. We have shown that LT-ADMM converges to a neighborhood of a stationary point, while LT-ADMM-VR converges exactly. We have thoroughly compared our algorithms with the state of the art, both theoretically and in simulations. Future research will focus on analyzing convergence for strongly convex problems, and extending our algorithm design to asynchronous scenarios.
Appendix A Preliminary Analysis
In this section we summarize the step-size bounds, and present preliminary results underpinning Theorems 1 and 2.
A-A Step-size bounds
The step-size upper bounds for LT-ADMM and LT-ADMM-VR are, respectively:
$\bar{\gamma}_{\text{sgd}}\coloneqq\min_{i=1,2,\ldots,7}\bar{\gamma}_{i},$ (10)
$\bar{\gamma}_{\text{sarah}}\coloneqq\min_{i=1,8,9,\ldots,17}\bar{\gamma}_{i},$ (11)
where:
$\bar{\gamma}_{1}$ $\coloneqq\min\left\{1,\frac{2}{|\tilde{\lambda}_{\min}|\rho\tau}\right\},\ \bar{\gamma}_{2}\coloneqq\frac{1}{4\sqrt{\tau(\tau-1)(L^{2}+\rho^{2}d_{u}^{2})}},$
$\bar{\gamma}_{3}$ $\coloneqq{\frac{3}{16L\tau}},\quad\bar{\gamma}_{4}\coloneqq\frac{1}{8\tau\sqrt{L^{2}+4L\rho^{2}d_{u}^{2}\tau+4\rho^{2}d_{u}^{2}}},$
$\bar{\gamma}_{5}$ $\coloneqq\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}}{4(\beta_{0}\kappa_{1}+\beta_{1})},\bar{\gamma}_{6}\coloneqq\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}}{8(\beta_{0}\kappa_{1}+\beta_{1})},$
$\bar{\gamma}_{7}$ $\coloneqq\max\left\{\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}}{256\kappa_{4}\tau^{2}N\beta_{0}},\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}\tau^{2}}{48L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}N}\right\}$
$\bar{\gamma}_{8}$ $\coloneqq\frac{m_{l}}{8L^{2}m_{u}},\quad\bar{\gamma}_{9}\coloneqq\min\left\{\frac{1}{2(4L^{2}+2\rho^{2}d_{u}^{2})},\frac{1}{2}\right\}$
$\bar{\gamma}_{10}$ $\coloneqq\frac{\sqrt{m_{l}}}{\tau L\sqrt{384m_{u}}},\quad\bar{\gamma}_{11}\coloneqq\frac{1}{2\sqrt{6\tau(\tau-1)(3L^{2}+\rho^{2}d_{u}^{2})}}$
$\bar{\gamma}_{12}$ $\coloneqq\frac{1}{8\sqrt{2\alpha_{0}\tau^{2}N+(12\alpha_{0}\tau L^{2}+\frac{L^{3}}{N})\tilde{s}_{2}}},\quad\bar{\gamma}_{13}\coloneqq{\frac{3}{16L\tau}}$
$\bar{\gamma}_{14}$ $\coloneqq\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}}{4(\tilde{\beta}_{0}\kappa_{1}+\tilde{\beta}_{1}+(96\tau^{2}L^{2}\tilde{\beta}_{0}+2\tilde{\beta}_{3})(\tilde{s}_{0}+\tilde{s}_{1}))}$
$\bar{\gamma}_{15}$ $\coloneqq\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}}{8(\tilde{\beta}_{0}\kappa_{1}+\tilde{\beta}_{1}+(96\tau^{2}L^{2}\tilde{\beta}_{0}+2\tilde{\beta}_{3})(\tilde{s}_{0}+\tilde{s}_{1}))}$
$\bar{\gamma}_{16}$ $\coloneqq\frac{24L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}N}{32\tau^{2}N\tilde{\beta}_{0}+192\tilde{s}_{2}\tau^{2}L^{2}\tilde{\beta}_{0}+4\tilde{s}_{2}\tilde{\beta}_{3}}$
$\bar{\gamma}_{17}$ $\coloneqq\frac{|\tilde{\lambda}_{\max}|^{2}\rho}{48\mu L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}N}.$
These bounds depend on the following quantities. $d_{u}=\max\{|\mathcal{N}_{i}|\}_{{i\in\mathcal{V}}}$ denotes the maximum agents’ degree. $\widehat{\mathbf{V}}$ is defined in (25). $\tilde{\lambda}_{\min}=\lambda_{\min}(\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}})$ and $\tilde{\lambda}_{\max}=\lambda_{\max}(\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}})$, with $-\tilde{\mathbf{L}}$ being the graph Laplacian (see (15)), and $\widehat{\mathbf{Q}}$ is being defined at the beginning of Appendix A-C. We denote $m_{u}=\max_{i=1,...,N}m_{i}$ and $m_{l}=\min_{i=1,...,N}m_{i}$, where $m_{i}$ is the number of local data points of agent $i$. Additionally, we have the following definitions, used both in the upper bound above and throughout the convergence analysis:
$\beta_{0}$ $\coloneqq 4(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}(L^{2}+\rho^{2}d_{u}^{2})$
$+{12L^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}\rho^{2}d_{u}^{2}$
$\beta_{1}$ $\coloneqq(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}\tau\rho d_{u}^{2}{72}/{|\tilde{\lambda}_{\max}|}$
$+L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}18\rho d_{u}^{2}\tau{216}/{|\tilde{\lambda}_{\max}|}$
$\kappa_{1}$ $\coloneqq\frac{72\tau}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{2}$
$\kappa_{4}$ $\coloneqq\frac{8}{N}\left((\frac{72\tau}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{2})\right.$
×$\left.\times(\frac{L^{2}}{2}+{2L\rho^{2}d_{u}^{2}\tau+2\rho^{2}d_{u}^{2}})\right.$
$\left.+\frac{2\tau}{|\tilde{\lambda}_{\max}|\rho}(18\rho^{2}d_{u}^{2}+18\rho^{2}d_{u}^{2}\tau L)\right),$
$\tilde{s}_{0}$ $\coloneqq\frac{36m_{u}}{|\tilde{\lambda}_{\max}|\rho}+\frac{16m_{u}}{m_{l}}\left(18+\frac{36\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)$
$\tilde{s}_{1}$ $\coloneqq\left(\frac{72\tau}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{2}\right)\frac{4m_{u}}{m_{l}}$
$\tilde{s}_{2}$ $\coloneqq\frac{64m_{u}}{m_{l}}\tau^{2}N+\frac{16m_{u}}{m_{l}}N,$
$\tilde{\beta}_{0}$ $\coloneqq(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}4(3L^{2}+\rho^{2}d_{u}^{2})$
$+{6L^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}(2\rho^{2}d_{u}^{2}+4L^{2})$
$\tilde{\beta}_{1}$ $\coloneqq(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}\tau\rho d_{u}^{2}{72}/{|\tilde{\lambda}_{\max}|}$
$+L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}18\rho d_{u}^{2}\tau{216}/{|\tilde{\lambda}_{\max}|}$
$\tilde{\beta}_{3}$ $\coloneqq(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}8L^{2}+{6L^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}4L^{2},$
$\alpha_{0}$ $\coloneqq\frac{L^{2}}{2N}+\frac{2L\rho^{2}d_{u}^{2}\tau+2\rho^{2}d_{u}^{2}+4\tau L^{3}}{N}$
$\alpha_{1}$ $\coloneqq\frac{72\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}$
$\alpha_{5}$ $\coloneqq\left({36\rho d_{u}^{2}\tau^{2}+36\rho d_{u}^{2}\tau^{3}L}\right)/(N|\tilde{\lambda}_{\max}|)$
$\mu$ $\coloneqq 32(\alpha_{5}+\alpha_{0}\alpha_{1}+2(48\tau^{2}L^{2}\alpha_{0}+{4\tau L^{3}}/{N})(\tilde{s}_{0}+\tilde{s}_{1})).$
A-B Preliminary transformation
We start by rewriting the algorithm in a compact form. To this end, we introduce the following auxiliary variables: $\mathbf{Z}=\operatorname{col}\{z_{ij}\}_{i,j\in\mathcal{E}}$, $\mathbf{\Phi}_{k}^{t}=\operatorname{col}\{\phi^{t}_{1,k},\phi_{2,k}^{t},...,\phi_{N,k}^{t}\}$, $G(\Phi_{k}^{t})=\operatorname{col}\{g_{1}(\phi_{1,k}^{t}),g_{2}(\phi_{2,k}^{t}),...,g_{N}(\phi_{N,k}^{t})\}$, $\mathrm{F}(\mathbf{X})=\operatorname{col}\{\frac{1}{N}f_{1}(x_{1}),\frac{1}{N}f_{2}(x_{2}),...,\frac{1}{N}f_{N}(x_{N})\}$. Define $\mathbf{A}=\operatorname{blk\,diag}\{\mathbf{1}_{d_{i}}\}_{i\in\mathcal{V}}\otimes\mathbf{I}_{n}\in\mathbb{R}^{Mn\times Nn},$ where $d_{i}=|\mathcal{N}_{i}|$ is the degree of node $i$, and $M=\sum_{i}|\mathcal{N}_{i}|$. $\mathbf{P}\in\mathbb{R}^{Mn\times Mn}$ is a permutation matrix that swaps $e_{ij}$ with $e_{ji}$. If there is an edge between nodes $i$, $j$, then $A^{T}[i,:]PA[:,j]=1$, otherwise $A^{T}[i,:]PA[:,j]=0$. Therefore $\mathbf{A}^{T}\mathbf{P}\mathbf{A}=\tilde{\mathbf{A}}$ is the adjacency matrix.
The compact form of LT-ADMM and LT-ADMM-VR then is:
$\mathbf{X}_{k+1}=\mathbf{X}_{k}-\gamma\sum_{t=0}^{\tau-1}(\nabla G(\mathbf{\Phi}_{k}^{t})+\rho\mathbf{A}^{T}\mathbf{A}\mathbf{\Phi}_{k}^{t}-\mathbf{A}^{T}\mathbf{Z}_{k})$ (12a)
$\mathbf{Z}_{k+1}=\frac{1}{2}\mathbf{Z}_{k}-\frac{1}{2}\mathbf{P}\mathbf{Z}_{k}+\rho\mathbf{P}\mathbf{A}\mathbf{X}_{k+1}.$ (12b)
Moreover, we introduce the following useful variables
$\mathbf{Y}_{k}=\mathbf{A}^{T}\mathbf{Z}_{k}-\nabla F(\mathbf{\bar{X}}_{k})-\rho\mathbf{D}\mathbf{X}_{k}$ (13)
$\tilde{\mathbf{Y}}_{k}=\mathbf{A}^{T}\mathbf{P}\mathbf{Z}_{k}+\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})-\rho\mathbf{D}\mathbf{X}_{k},$
where $\bar{\mathbf{X}}_{k}=\mathbf{1}_{N}\otimes\bar{x}_{k}$, with $\bar{x}_{k}=\frac{1}{N}\mathbf{1}^{T}\mathbf{X}_{k}$, and $\mathbf{D}=\mathbf{A}^{T}\mathbf{A}=\operatorname{diag}\{d_{i}\mathbf{I}_{n}\}_{i\in\mathcal{V}}$ is the degree matrix.
Multiplying both sides of (12b) by $\mathbf{1}^{T}$, and using the initial condition, we obtain $\mathbf{1}^{T}\mathbf{A}^{T}\mathbf{Z}_{k+1}=\rho\mathbf{1}^{T}\mathbf{D}\mathbf{X}_{k+1}$ for all $k\in\mathbb{N}$. As a consequence $\bar{\mathbf{Y}}_{k}=\mathbf{1}\otimes\frac{1}{N}\mathbf{1}^{T}\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})=\mathbf{1}\otimes\frac{1}{N}\sum_{i}\nabla f_{i}(\bar{x}_{k})$, and (12) can be further rewritten as
$\begin{bmatrix}\mathbf{X}_{k+1}\\ \mathbf{Y}_{k+1}\\ \tilde{\mathbf{Y}}_{k+1}\end{bmatrix}=\begin{bmatrix}\mathbf{I}&\gamma\tau\mathbf{I}&\mathbf{0}\\ \rho\tilde{\mathbf{L}}&\rho\tilde{\mathbf{L}}\gamma\tau+\frac{1}{2}\mathbf{I}&-\frac{1}{2}\mathbf{I}\\ \mathbf{0}&-\frac{1}{2}\mathbf{I}&\frac{1}{2}\mathbf{I}\end{bmatrix}\otimes\mathbf{I}_{n}\begin{bmatrix}\mathbf{X}_{k}\\ \mathbf{Y}_{k}\\ \tilde{\mathbf{Y}}_{k}\end{bmatrix}-\mathbf{h}_{k},$ (14)
where
$\tilde{\mathbf{L}}=\tilde{\mathbf{A}}-\mathbf{D}$ (15)
and
$\mathbf{h}_{k}$ $=\large[\gamma\sum_{t=0}^{\tau-1}(\nabla G(\mathbf{\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho D\mathbf{\Phi}_{k}^{t}-\rho\mathbf{D}{\mathbf{X}}_{k});$
$\gamma\rho\tilde{\mathbf{L}}\sum_{t=0}^{\tau-1}(\nabla G(\mathbf{\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\mathbf{\Phi}_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})$
$+\nabla F(\bar{\mathbf{X}}_{k+1})-\nabla F(\bar{\mathbf{X}}_{k});-\nabla F(\bar{\mathbf{X}}_{k+1})+\nabla F(\bar{\mathbf{X}}_{k})\large].$
We remark that (14) can be interpreted as a linear dynamical system, with the non-linearity of the gradients as input in $\mathbf{h}_{k}$.
A-C Deviation from the average
The following lemma illustrates how far the states deviate from the average and will be used later in the proofs of Lemmas 2 and 6.
Lemma 1.
Let Assumption 1 hold, when $\gamma<\frac{2}{\tau|\tilde{\lambda}_{\min}|\rho}$,
$\|\bar{\mathbf{X}}_{k}-\mathbf{X}_{k}\|^{2}\leq\frac{18\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}\|\widehat{\mathbf{d}}_{k}\|^{2},\quad\|\bar{\mathbf{Y}}_{k}-\mathbf{Y}_{k}\|^{2}\leq 9\|\widehat{\mathbf{d}}_{k}\|^{2},$ (16)
and
$\|\widehat{\mathbf{d}}_{k+1}\|^{2}\leq\delta\|\widehat{\mathbf{d}}_{k}\|^{2}+\frac{1}{1-\delta}\|\mathbf{\widehat{h}}_{k}\|^{2}$ (17)
where $\delta=1-{|\tilde{\lambda}_{\max}|\rho\tau\gamma}/{2}<1$.
Proof.
By Assumption 1, graph $\mathcal{G}$ is undirected and connected, hence its Laplacian $-\tilde{\mathbf{L}}$ is symmetric; moreover, it has one zero eigenvalue with eigenvector $\boldsymbol{1}$, with all eigenvalues being positive. Denote by $\widehat{\mathbf{Q}}\in\mathbf{R}^{N\times(N-1)}$ the matrix satisfying $\widehat{\mathbf{Q}}\widehat{\mathbf{Q}}^{T}=\mathbf{I}_{N}-\frac{1}{N}\mathbf{11}{}^{T}$, $\widehat{\mathbf{Q}}^{T}\widehat{\mathbf{Q}}=\mathbf{I}_{N-1}$ and $\mathbf{1}^{T}\widehat{\mathbf{Q}}=0$, $\widehat{\mathbf{Q}}^{T}\mathbf{1}=0$. We have that
$\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}=\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}(\mathbf{I}_{N}-\frac{1}{N}\mathbf{11}{}^{T})=\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}\widehat{\mathbf{Q}}^{T}.$ (18)
Additionally, it holds that $\|\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k}\|^{2}=\mathbf{X}_{k}^{T}\widehat{\mathbf{Q}}\widehat{\mathbf{Q}}^{T}\widehat{\mathbf{Q}}\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k}=\left\|\widehat{\mathbf{Q}}\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k}\right\|^{2}=\left\|\mathbf{X}_{k}-\bar{{\mathbf{X}}}_{k}\right\|^{2}$, and $\|\widehat{\mathbf{Q}}\|=1$. Multiplying both sides of (14) by $\widehat{\mathbf{Q}}^{T}$ and using (18) yields:
$\begin{bmatrix}\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k+1}\\ \widehat{\mathbf{Q}}^{T}\mathbf{Y}_{k+1}\\ \widehat{\mathbf{Q}}^{T}\tilde{\mathbf{Y}}_{k+1}\end{bmatrix}$ $=(\mathbf{\Theta}\otimes\mathbf{I}_{n})\begin{bmatrix}\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k}\\ \widehat{\mathbf{Q}}^{T}\mathbf{Y}_{k}\\ \widehat{\mathbf{Q}}^{T}\tilde{\mathbf{Y}}_{k}\end{bmatrix}-\widehat{\mathbf{Q}}^{T}\mathbf{h}_{k}$ (19)
where $\mathbf{\Theta}=\begin{bmatrix}\mathbf{I}&\gamma\tau\mathbf{I}&\mathbf{0}\\ \rho\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}&\rho\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}\gamma\tau\mathbf{I}+\frac{1}{2}\mathbf{I}&-\frac{1}{2}\mathbf{I}\\ \mathbf{0}&\frac{1}{2}\mathbf{I}&\frac{1}{2}\mathbf{I}\end{bmatrix}$.
The next step is to show that $\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}$ is negative definite by contradiction. Let $x\in R^{N-1}$ be an arbitrary vector, since $-\tilde{\mathbf{L}}$ is the positive semi-definite Laplacian matrix, the quadratic form $x^{T}\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}x=(\widehat{\mathbf{Q}}x)^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}x\leq 0$. Moreover, if $(\widehat{\mathbf{Q}}x)^{T}(\tilde{\mathbf{A}}-\mathbf{D})\widehat{\mathbf{Q}}x=0$, we have $\widehat{\mathbf{Q}}x=\boldsymbol{1}$. Now, the properties of $\widehat{\mathbf{Q}}$ imply that $\widehat{\mathbf{Q}}^{T}\widehat{\mathbf{Q}}x=x=\widehat{\mathbf{Q}}^{T}\boldsymbol{1}=0$. Therefore, for all non-zero vectors $x$, the quadratic form $x^{T}\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}x<0$, thus $\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}$ is a symmetric negative-definite matrix.
We proceed now to diagonalize each block of $\mathbf{\Theta}$ with $\boldsymbol{\phi}\in\mathbb{R}^{(N-1)\times(N-1)}$:
$\tilde{\mathbf{\Theta}}$ $=\boldsymbol{\phi}\mathbf{\Theta}\boldsymbol{\phi}^{T}=\begin{bmatrix}\phi&0&0\\ 0&\phi&0\\ 0&0&\phi\\ \end{bmatrix}\mathbf{\Theta}\begin{bmatrix}\phi^{T}&0&0\\ 0&\phi^{T}&0\\ 0&0&\phi^{T}\\ \end{bmatrix}$
$=\begin{bmatrix}\mathbf{I}&\gamma\tau&0\\ \rho\phi\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}\phi^{T}&\rho\phi\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}\phi^{T}\gamma\tau+\frac{1}{2}\mathbf{I}&-\frac{1}{2}\mathbf{I}\\ 0&-\frac{1}{2}\mathbf{I}&\frac{1}{2}\mathbf{I}\end{bmatrix}.$
We denote $\phi\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}\phi^{T}=\operatorname{diag}\{\tilde{\lambda}_{i}\}_{i=2,...,N}$, where $\tilde{\lambda}_{i}<0$ is the eigenvalue of $\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}}$, $\tilde{\lambda}_{\min}=\lambda_{\min}(\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}})$, and $\tilde{\lambda}_{\max}=\lambda_{\max}(\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{L}}\widehat{\mathbf{Q}})$. Since each block of $\tilde{\mathbf{\Theta}}$ is a diagonal matrix, there exists a permutation matrix $\mathbf{P}_{0}$ such that $\mathbf{P}_{0}\tilde{\mathbf{\Theta}}\mathbf{P}_{0}^{T}=\mathbf{P}_{0}\boldsymbol{\phi}\mathbf{\Theta}\boldsymbol{\phi}^{T}\mathbf{P}_{0}^{T}=\operatorname{blkdiag}\left\{\mathbf{D}_{i}\right\}_{i=2}^{N},$ where
$\mathbf{D}_{i}=\begin{bmatrix}1&\gamma\tau&0\\ \rho\tilde{\lambda}_{i}&\rho\tilde{\lambda}_{i}\gamma\tau+0.5&-0.5\\ 0&-0.5&0.5\end{bmatrix}.$ (20)
We diagonalize $\mathbf{D}_{i}=\mathbf{V}_{i}\mathbf{\Delta}_{i}\mathbf{V}_{i}^{-1}$, where $\mathbf{\Delta}_{i}$ is the diagonal matrix of $\mathbf{D}_{i}$’s eigenvalues, and
$\mathbf{V}_{i}=\begin{bmatrix}-\gamma\tau&d_{12}&d_{13}\\ 1&d_{22}&d_{23}\\ 1&1&1\end{bmatrix}$ (21)
with $d_{12}=-\gamma\tau+((\gamma\tilde{\lambda_{i}}\rho\tau(\gamma\tilde{\lambda_{i}}\rho\tau+2))^{0.5})/(\tilde{\lambda_{i}}\rho)$, $d_{13}=-\gamma\tau-((\gamma\tilde{\lambda_{i}}\rho\tau(\gamma\tilde{\lambda_{i}}\rho\tau+2))^{0.5})/(\tilde{\lambda_{i}}\rho)$, $d_{22}=\tilde{\lambda_{i}}\rho d_{12}-1$, $d_{23}=\tilde{\lambda_{i}}\rho d_{13}-1$. The nonzero eigenvalues $\lambda$ of $\mathbf{D}_{i}$, $i=2,\ldots,N$, satisfy $2\lambda^{2}+(-2\tilde{\lambda}_{i}\rho\tau\gamma-4)\lambda+\tilde{\lambda}_{i}\rho\tau\gamma+2=0,$ which can be written in the form:
$02\lambda^{2}-2t\lambda+t=$ 0 (22)
where $t=\tilde{\lambda}_{i}\rho\tau\gamma+2$. The modulus of the roots of (22) is $1-\frac{|\tilde{\lambda}|\rho\tau\gamma}{2}$ when $-2<\tilde{\lambda}\rho\tau\gamma<0$. We conclude that we can write $\mathbf{\Theta}=(\mathbf{P}_{0}\boldsymbol{\phi})^{T}\mathbf{V}\boldsymbol{\Delta}\mathbf{V}^{-1}(\mathbf{P}_{0}\boldsymbol{\phi})$ where $\mathbf{V}=\operatorname{blkdiag}\left\{V_{i}\right\}_{i=2}^{N}$ and
$\boldsymbol{\Delta}=\operatorname{blkdiag}\left\{\mathbf{\Delta}_{i}\right\}_{i=2}^{N}.$ (23)
Moreover, $\|\mathbf{\Delta}\|=1-{|\tilde{\lambda}_{\max}|\rho\tau\gamma}/{2}$ when
$|\tilde{\lambda}_{\min}|\rho\tau\gamma<2.$ (24)
Then, left multiplying both sides of (19) by the inverse of
$\widehat{\mathbf{V}}=(P_{0}\boldsymbol{\phi})^{T}\mathbf{V},$ (25)
which is given by $\widehat{\mathbf{V}}^{-1}=\mathbf{V}^{-1}(\mathbf{P}_{0}\boldsymbol{\phi})$, yields
$\widehat{\mathbf{d}}_{k+1}=\mathbf{\Delta}\widehat{\mathbf{d}}_{k}-\widehat{\mathbf{h}}_{k},$ (26)
where $\widehat{\mathbf{d}}_{k}=\widehat{\mathbf{V}}^{-1}\begin{bmatrix}\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k};\widehat{\mathbf{Q}}^{T}\mathbf{Y}_{k};\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{Y}}_{k}\end{bmatrix}$, $\widehat{\mathbf{h}}_{k}=\widehat{\mathbf{V}}^{-1}\widehat{\mathbf{Q}}^{T}\mathbf{h}_{k}$, and $\begin{bmatrix}\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k};\widehat{\mathbf{Q}}^{T}\mathbf{Y}_{k};\widehat{\mathbf{Q}}^{T}\tilde{\mathbf{Y}}_{k}\end{bmatrix}=\widehat{\mathbf{V}}\widehat{\mathbf{d}}_{k}=\boldsymbol{\phi}^{T}\mathbf{P}_{0}^{T}\mathbf{V}\widehat{\mathbf{d}}_{k}=\boldsymbol{\phi}^{T}\mathbf{P}_{0}^{T}\mathbf{V}\mathbf{P}_{0}\mathbf{P}_{0}^{T}\widehat{\mathbf{d}}_{k}$. As a consequence, from (19) it holds that
$\begin{bmatrix}\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k}\\ \widehat{\mathbf{Q}}^{T}\mathbf{Y}_{k}\\ \widehat{\mathbf{Q}}^{T}\tilde{\mathbf{Y}}_{k}.\end{bmatrix}=\boldsymbol{\phi}^{T}\left[\begin{array}[]{ccc}-\gamma\tau\mathbf{I}&d_{12}\mathbf{I}&d_{13}\mathbf{I}\\ \mathbf{I}&d_{22}\mathbf{I}&d_{23}\mathbf{I}\\ \mathbf{I}&\mathbf{I}&\mathbf{I}\end{array}\right]\mathbf{P}_{0}^{T}\widehat{\mathbf{d}}_{k}$
$=\boldsymbol{\phi}^{T}\left[\begin{array}[]{c}-\gamma\tau\mathbf{I}\mathbf{P}_{0}^{T}[1]+d_{12}\mathbf{P}_{0}^{T}[2]+d_{13}\mathbf{P}_{0}^{T}[3]\\ \mathbf{P}_{0}^{T}[1]+d_{22}\mathbf{P}_{0}^{T}[2]+d_{23}\mathbf{P}_{0}^{T}[3]\\ \mathbf{P}_{0}^{T}[1]+\mathbf{P}_{0}^{T}[2]+\mathbf{P}_{0}^{T}[3]\end{array}\right]\widehat{\mathbf{d}}_{k},$
where $\mathbf{P}_{0}^{T}[1],\mathbf{P}_{0}^{T}[2],\mathbf{P}_{0}^{T}[3]$ are the top, middle and bottom blocks of $\mathbf{P}_{0}^{T}$ respectively. Moreover, we have $|d_{12}|^{2}=|d_{13}|^{2}\leq\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}$, $|d_{22}|=|d_{23}|=1$. Now, if we let $\gamma\tau\leq\frac{2}{|\tilde{\lambda}_{\max}|\rho},$ and using $\|\phi\|=1$, $\|\mathbf{P}_{0}^{T}[i]\|=1$, $i=1,2,3$, we derive that
$\|\bar{\mathbf{X}}_{k}-\mathbf{X}_{k}\|^{2}=\|\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k}\|^{2}$
$\quad=\|\phi^{T}(-\gamma\tau\mathbf{I}\mathbf{P}_{0}^{T}[1]+d_{12}\mathbf{P}_{0}^{T}[2]+d_{13}\mathbf{P}_{0}^{T}[3])\widehat{\mathbf{d}}_{k}\|^{2}$
$\quad\leq 3(\gamma^{2}\tau^{2}+\frac{4\gamma\tau}{|\tilde{\lambda}_{\max}|\rho})\|\widehat{\mathbf{d}}_{k}\|^{2}\leq\frac{18\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}\|\widehat{\mathbf{d}}_{k}\|^{2}.$
Applying the same manipulations to $\|\bar{\mathbf{Y}}_{k}-\mathbf{Y}_{k}\|^{2}$, we obtain (16) holds. Denote now $\|\widehat{\mathbf{\Phi}}_{k}\|^{2}=\sum_{i=1}^{N}\sum_{t=0}^{\tau-1}\left\|\phi_{i,k}^{t}-\bar{x}_{k}\right\|^{2}=\sum_{t=0}^{\tau-1}\left\|\Phi_{k}^{t}-\bar{X}_{k}\right\|^{2}.$ Using Assumption 2 we derive that
$\|\sum_{t=0}^{\tau-1}(G({\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}{\Phi}_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}$
$=\|\sum_{t=0}^{\tau-1}(G({\Phi}_{k}^{t})-\nabla F(\mathbf{\Phi}_{k}^{t})+\nabla F({\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})$
$+\rho\mathbf{D}({\Phi}_{k}^{t}-\bar{\mathbf{X}}_{k}+\bar{\mathbf{X}}_{k}-\mathbf{X}_{k}))\|^{2}$
$\leq 4(\tau L^{2}+\tau\rho^{2}d_{u}^{2})\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+4\tau^{2}\rho^{2}d_{u}^{2}\left\|\bar{\mathbf{X}}_{k}-\mathbf{X}_{k}\right\|^{2}$
$+4\tau\sum_{t=0}^{\tau-1}\|G({\Phi}_{k}^{t})-\nabla F({\Phi}_{k}^{t})\|^{2}.$
Denoting $\epsilon_{k}=\sum_{t=0}^{\tau-1}(\rho\mathbf{1}^{T}\mathbf{D}\Phi_{k}^{t}-\mathbf{1}^{T}\mathbf{A}^{T}\mathbf{Z}_{k})=\rho\sum_{t=0}^{\tau-1}\sum_{i=1}^{N}(d_{i}\phi_{i,k}^{t}-d_{i}\bar{x}_{k}+d_{i}\bar{x}_{k}-d_{i}x_{i,k})$, we have
$\epsilon_{k}^{2}\leq 2\rho^{2}d_{u}^{2}\tau N(\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\tau\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}).$ (27)
Now, denoting $\overline{G}(\boldsymbol{\Phi}^{t}_{k})=\frac{1}{N}\sum_{i}g_{i}(\phi_{i,k}^{t})$, then $\|\sum_{t}\overline{G}(\boldsymbol{\Phi}^{t}_{k})\|^{2}\leq 2\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}+\frac{2\tau}{N}\sum_{t}\sum_{i}\|\nabla f_{i}(\phi_{i,k}^{t})-g_{i}(\phi_{i,k}^{t})\|^{2},$ where $\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}=\|\frac{1}{N}\sum_{i}\sum_{t}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}\leq\frac{2\tau}{N}\sum_{i}\sum_{t}\|\nabla f_{i}(\phi_{i,k}^{t})-\nabla f_{i}(\bar{x}_{k})\|^{2}+2\tau^{2}\|\frac{1}{N}\sum_{i}\nabla f_{i}(\bar{x}_{k})\|^{2}.$ We also have
$\left\|\nabla F(\bar{\mathbf{X}}_{k+1})-\nabla F(\bar{\mathbf{X}}_{k})\right\|^{2}=N\left\|\nabla F(\bar{x}_{k+1})-\nabla F(\bar{x}_{k})\right\|^{2}$
$\leq NL^{2}\|\frac{\gamma}{N}\sum_{t=0}^{\tau-1}\sum_{i=1}^{N}g_{i}(\phi_{k}^{t,i})+\frac{\gamma}{N}\sum_{t=0}^{\tau-1}(\rho\mathbf{1}^{T}\mathbf{D}\Phi_{k}^{t}-\mathbf{1}^{T}\mathbf{A}^{T}\mathbf{Z}_{k})\|^{2}$
$\leq 2L^{2}\gamma^{2}\left(2\tau\sum_{t}\sum_{i}\|\nabla f_{i}\left(\phi_{i,k}^{t}\right)-g_{i}\left(\phi_{i,k}^{t}\right)\|^{2}\right.$
$\left.+2N\|\sum_{t}\overline{\nabla F}\left(\boldsymbol{\Phi}^{t}_{k}\right)\|^{2}+2\rho^{2}d_{u}^{2}\tau(\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\tau\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2})\right).$
Using $\|\widehat{\mathbf{Q}}^{T}\|=1$ it further holds that:
$\|\mathbf{\widehat{h}}_{k}\|^{2}\leq\gamma^{2}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}\cdot$
$\cdot\|\sum_{t=0}^{\tau-1}(G(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}$
$+3\|\widehat{\mathbf{V}}^{-1}\|^{2}\left\|\nabla F(\bar{\mathbf{X}}_{k+1})-\nabla F(\bar{\mathbf{X}}_{k})\right\|^{2}$
$\leq\gamma^{2}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}\left(4(\tau L^{2}+\tau\rho^{2}d_{u}^{2})\|\widehat{\mathbf{\Phi}}_{k}\|^{2}\right.$
$\left.+4\tau^{2}\rho^{2}d_{u}^{2}\left\|\bar{\mathbf{X}}_{k}-\mathbf{X}_{k}\right\|^{2}+4\tau\sum_{t=0}^{\tau-1}\|G(\Phi_{k}^{t})-\nabla F(\Phi_{k}^{t})\|^{2}\right)$
$+{6L^{2}\gamma^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}\cdot\left(2\tau\sum_{t}\sum_{i}\|\nabla f_{i}\left(\phi_{i,k}^{t}\right)-g_{i}\left(\phi_{i,k}^{t}\right)\|^{2}\right.$
$\left.+2N\|\sum_{t}\overline{\nabla F}\left({\Phi}^{t}_{k}\right)\|^{2}+2\rho^{2}d_{u}^{2}\tau(\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\tau\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2})\right).$ (28)
Recalling (26), using Jensen’s inequality $\|\widehat{\mathbf{d}}_{k+1}\|^{2}\leq\frac{1}{\|\mathbf{\Delta}\|}\|\mathbf{\Delta}\|^{2}\|\widehat{\mathbf{d}}_{k}\|^{2}+\frac{1}{1-\|\mathbf{\Delta}\|}\|\mathbf{\widehat{h}}_{k}\|^{2}$ yields (17). ∎
Appendix B Convergence analysis for LT-ADMM
B-A Key bounds
Lemma 2.
Let Assumptions 1, 2, and 4 hold; when $\gamma<\frac{2}{\tau|\tilde{\lambda}_{\max}|\rho}$ and $4\gamma^{2}\tau(L^{2}+\rho^{2}d_{u}^{2})<\frac{1/4}{\tau-1}$, we have
$\mathbb{E}\left[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}\right]$ (29)
$\leq\left(\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2}\right)\mathbb{E}\left[\|\widehat{\mathbf{d}}_{k}\|^{2}\right]$
$+16\tau^{3}\gamma^{2}N\mathbb{E}\left\|\nabla F(\bar{x}_{k})\right\|^{2}+4N\tau^{2}\gamma^{2}\sigma^{2}.$
Proof.
From (12) we can derive that
$\bar{x}_{k+1}-x^{*}$ $=\bar{x}_{k}-x^{*}-\frac{\gamma}{N}\sum_{t=0}^{\tau-1}\sum_{i=1}^{N}g_{i}(\phi_{i,k}^{t})$ (30)
$-\frac{\gamma}{N}\left(\sum_{t=0}^{\tau-1}(\rho\mathbf{1}^{T}\mathbf{D}\Phi_{k}^{t}-\mathbf{1}^{T}\mathbf{A}^{T}\mathbf{Z}_{k})\right)$
and
$\Phi_{k}^{t+1}=\Phi_{k}^{t}+\gamma\mathbf{Y}_{k}-\gamma(G(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})$ (31)
Recall that by Assumption 4, $\|G(\Phi_{k}^{t})-\nabla F(\Phi_{k}^{t})\|^{2}\leq N\sigma^{2}$.
Now, suppose that $\tau\geq 2$, using Jensen’s inequality we obtain
$\mathbb{E}[\left\|\Phi_{k}^{t+1}-\bar{\mathbf{X}}_{k}\right\|^{2}]$
$\leq\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}+\gamma\mathbf{Y}_{k}-\gamma(\nabla F(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+$
$+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}]+N\gamma^{2}\sigma^{2}$
$\leq(1+\frac{1}{\tau-1})\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}]+N\gamma^{2}\sigma^{2}+$
$+\tau\gamma^{2}\mathbb{E}[\|\mathbf{Y}_{k}-(\nabla F(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}]$
${\leq}(1+\frac{1}{\tau-1})\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}]+N\gamma^{2}\sigma^{2}+$
$+2\tau\gamma^{2}\mathbb{E}[\|\nabla F(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\bar{\mathbf{X}}_{k}\|^{2}]+$
$+2\tau\gamma^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}]$
$\leq(1+\frac{1}{\tau-1}+4\gamma^{2}\tau(L^{2}+\rho^{2}d_{u}^{2}))\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}+$
$+2\tau\gamma^{2}\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}]+N\gamma^{2}\sigma^{2}$
$\leq(1+\frac{5/4}{\tau-1})\mathbb{E}[\left\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}]+N\gamma^{2}\sigma^{2}+$
$+2\tau\gamma^{2}\mathbb{E}[\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}],$
where $\bar{\mathbf{Y}}_{k}=\mathbf{1}\otimes\frac{1}{N}\mathbf{1}^{T}\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})=\mathbf{1}\otimes\frac{1}{N}\sum_{i}\nabla f_{i}(\bar{x}_{k})$, and the last inequality holds when
$4\gamma^{2}\tau(L^{2}+\rho^{2}d_{u}^{2})<\frac{1/4}{\tau-1}.$ (32)
Iterating the above inequality for $t=0,...,\tau-1$
$\mathbb{E}$ $[\left\|\Phi_{k}^{t+1}-\bar{\mathbf{X}}_{k}\right\|^{2}]\leq\left(1+\frac{5/4}{\tau-1}\right)^{t}\mathbb{E}[\left\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\right\|^{2}]+$
$+2\tau\gamma^{2}\sum_{l=0}^{t}\left(1+\frac{5/4}{\tau-1}\right)^{l}\cdot$
$\cdot\mathbb{E}[\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}]+$
$+N\gamma^{2}\sigma^{2}\sum_{l=0}^{t}\left(1+\frac{5/4}{\tau-1}\right)^{l}$
$\leq 4\mathbb{E}[\left\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\right\|^{2}]+4\tau N\gamma^{2}\sigma^{2}+$
$+8\tau^{2}\gamma^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}],$
where the last inequality holds by $(1+\frac{a}{\tau-1})^{t}\leq\exp(\frac{at}{\tau-1})\leq\exp(a)$ for $t\leq\tau-1$ and $a={5}/{4}$.
Summing over $t$, it follows that
$\begin{split}&\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]\leq 4\tau\mathbb{E}[\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}]+16\tau^{3}\gamma^{2}N\mathbb{E}[\left\|\nabla F(\bar{x}_{k})\right\|^{2}]+\\ &+4N\tau^{2}\gamma^{2}\sigma^{2}+16\tau^{3}\gamma^{2}\mathbb{E}[\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}];\end{split}$ (33)
moreover, it is easy to verify that (33) also holds for $\tau=1$. As a consequence of (16), we obtain
$\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}\leq(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})\|\widehat{\mathbf{d}}_{k}\|^{2},$ (34)
and using this fact into (33) concludes the proof. ∎
Lemma 3.
Let Assumptions 1, 2, and 4 hold; when $\gamma\leq\min\{\bar{\gamma}_{1},\bar{\gamma}_{2}\}$, it holds that
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k+1}\|^{2}]\leq(\delta+\frac{q_{1}}{1-\delta})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+$ (35)
$+\frac{q_{2}}{1-\delta}\mathbb{E}[\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}]+\frac{q_{3}}{1-\delta}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]+q_{4}\sigma^{2}$
where
$\delta$ $\coloneqq 1-\frac{|\tilde{\lambda}_{\max}|\rho\tau\gamma}{2},q_{4}\coloneqq\frac{4N\tau^{2}\gamma^{2}c_{0}+c_{3}}{1-\delta}$
$c_{0}$ $\coloneqq\gamma^{2}\tau\beta_{0},c_{1}\coloneqq\gamma^{3}\tau^{2}\beta_{1},c_{2}\coloneqq{12L^{2}\gamma^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}N,$
$c_{3}$ $\coloneqq 4\gamma^{2}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}\tau^{2}N+{12L^{2}\gamma^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}\tau^{2}N,$
$q_{1}$ $\coloneqq c_{0}\kappa_{1}\gamma\tau+c_{1},q_{2}\coloneqq c_{2},q_{3}\coloneqq 16\tau^{3}\gamma^{2}Nc_{0}.$
Proof.
When $\gamma\leq\min\{\bar{\gamma}_{1},\bar{\gamma}_{2}\}$, it follows that
$\gamma\tau\leq\frac{2}{|\tilde{\lambda}_{\max}|\rho},\quad|\tilde{\lambda}_{\min}|\rho\tau\gamma<2,\quad\gamma^{2}\tau(L^{2}+\rho^{2}d_{u}^{2})<\frac{1/16}{\tau-1}.$ (36)
Using (28) and Assumption 4, we have
$\|\widehat{\mathbf{h}}_{k}\|^{2}\leq c_{0}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+c_{1}\|\widehat{\mathbf{d}}_{k}\|^{2}+c_{2}\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}+c_{3}\sigma^{2};$
using this fact together with (29) and (17), and by $\gamma<1$, we can then derive that (35) holds. ∎
B-B Theorem 1
We start our proof by recalling that the following inequality holds for and $L$-smooth function $f$, $\forall y,z\in\mathbb{R}^{n}$ [36]:
$f(y)\leq f(z)+\langle\nabla f(z),y-z\rangle+(L/2)\|y-z\|^{2}$ (37)
Based on (30), substituting $y=\bar{x}_{k+1}$ and $z=\bar{x}_{k}$ into (37), using Assumption 4 and (27), we get
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]$
$\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\gamma\mathbb{E}[\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}(\sum_{t}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)+\epsilon_{k})\rangle]$
$+\frac{\gamma^{2}L}{2}\mathbb{E}[\|\frac{1}{N}(\sum_{t}\sum_{i}g_{i}(\phi_{i,k}^{t})+\epsilon_{k})\|^{2}]$
$\leq\mathbb{E}[F(\bar{x}_{k})]-\gamma[\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}(\sum_{t}\sum_{i}\nabla f_{i}(\phi_{i,k}^{t})+\epsilon_{k})\rangle]$
$+2\gamma^{2}\tau L\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]+{2\gamma^{2}\tau^{2}L\sigma^{2}}$
$+\frac{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau\mathbb{E}\left[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\tau\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}\right]}{N}.$
Using now $2\langle a,b\rangle=\|a\|^{2}+\|b\|^{2}-\|a-b\|^{2}$, we have
$-\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}\sum_{t}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)+\frac{1}{N}\epsilon_{k}\rangle$
$\leq-\sum_{t}\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}\sum_{i}\nabla f_{i}(\phi_{i,k}^{t})\rangle+\frac{\tau}{4}\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}$
$+\frac{1}{N^{2}\tau}\epsilon_{k}^{2}$
$=-\frac{\tau}{2}\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}-\frac{1}{2}\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}$
$+\frac{1}{2}\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)-\nabla F\left(\bar{x}_{k}\right)\|^{2}$
$+\frac{\tau}{4}\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}+\frac{1}{N^{2}\tau}\epsilon_{k}^{2}$
$\leq-\frac{\tau}{4}\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}-\frac{1}{2}\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}$
$+(\frac{L^{2}}{2N}+\frac{2\rho^{2}d_{u}^{2}}{N})\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\frac{2\rho^{2}d_{u}^{2}\tau}{N}\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}$
Now, combining the two equations above, and using (16), yields
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\frac{\gamma\tau}{4}\mathbb{E}[\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}]$
$-\frac{\gamma}{2}(1-4\gamma\tau L)\sum_{t}\mathbb{E}[\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\left(\frac{\gamma L^{2}}{2N}+\frac{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}}{N}\right)\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]$
$+\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}\frac{18\gamma\rho^{2}d_{u}^{2}\tau+18\gamma^{2}\rho^{2}d_{u}^{2}\tau^{2}L}{N}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+{2\gamma^{2}\tau^{2}L\sigma^{2}}.$
Substituting (29) into the above inequality yields
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]+$
$-\frac{\gamma\tau}{4}\left(1-64\gamma\tau^{2}(\frac{\gamma L^{2}}{2}+2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2})\right)\cdot$
$\cdot\mathbb{E}[\|\nabla F\left(\bar{x}_{k}\right)\|^{2}]$
$-\frac{\gamma}{2}(1-4\gamma L\tau)\sum_{t}\mathbb{E}[\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\frac{1}{N}\left((\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2})\cdot\right.$
$\left.\cdot(\frac{\gamma L^{2}}{2}+{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}})\right.$
$\left.+\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}({18\gamma\rho^{2}d_{u}^{2}\tau+18\gamma^{2}\rho^{2}d_{u}^{2}\tau^{2}L})\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+{2\gamma^{2}\tau^{2}L\sigma^{2}}+\left(\frac{\gamma L^{2}}{2}+{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}}\right)4\tau^{2}\gamma^{2}\sigma^{2}.$
When $\gamma\leq\min\{\bar{\gamma}_{3},\bar{\gamma}_{4}\}$, then
$\gamma\tau^{2}(\frac{\gamma L^{2}}{2}+2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2})<\frac{1}{128},\quad\gamma L\tau<\frac{3}{16},$ (38)
and we can upper bound the previous inequality by
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\frac{\gamma\tau}{8}\mathbb{E}[\|\nabla F\left(\bar{x}_{k}\right)\|^{2}]+$
$-\frac{\gamma}{8}\sum_{t}\mathbb{E}[\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\frac{1}{N}\left((\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2})\cdot\right.$
$\left.\cdot(\frac{\gamma L^{2}}{2}+{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}})\right.$
$\left.+\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}({18\gamma\rho^{2}d_{u}^{2}\tau+18\gamma^{2}\rho^{2}d_{u}^{2}\tau^{2}L})\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+{2\gamma^{2}\tau^{2}L\sigma^{2}}+\left(\frac{\gamma L^{2}}{2}+{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}}\right)4\tau^{2}\gamma^{2}\sigma^{2}.$
Rearranging the above relation we get
$\mathcal{D}_{k}\leq\frac{8}{\gamma\tau}\mathbb{E}\left[\left(\tilde{F}\left(\bar{x}_{k}\right)-\tilde{F}\left(\bar{x}_{k+1}\right)\right)\right]$
$+\frac{8}{N\gamma\tau}\left((\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2})\cdot\right.$
$\left.\cdot(\frac{\gamma L^{2}}{2}+{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}})\right.$
$\left.+\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}({18\gamma\rho^{2}d_{u}^{2}\tau+18\gamma^{2}\rho^{2}d_{u}^{2}\tau^{2}L})\right)\cdot$
$\cdot\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+16\left({\gamma L^{2}}+{4\gamma^{2}L\rho^{2}d_{u}^{2}\tau+4\gamma\rho^{2}d_{u}^{2}}\right)\tau\gamma\sigma^{2}$
$+{16\gamma\tau L\sigma^{2}},$
where $\mathcal{D}_{k}$ is defined in (7), and $\tilde{F}\left(\bar{x}_{k}\right)=F\left(\bar{x}_{k}\right)-F(x^{*})$.
Summing over $k=0,1,\ldots,K-1$, using $-\tilde{F}\left(\bar{x}_{k}\right)\leq 0$ and $\gamma\leq 1$, it holds that
$\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\frac{8\tilde{F}\left(\bar{x}^{0}\right)}{\gamma\tau}+c_{4}\sum_{k=0}^{K-1}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+Kc_{5}\sigma^{2}$ (39)
where $c_{4}\coloneqq\gamma\kappa_{4}$, $c_{5}\coloneqq{16\gamma\tau L}+16\left({\gamma L^{2}}+{4\gamma^{2}L\rho^{2}d_{u}^{2}\tau+4\gamma\rho^{2}d_{u}^{2}}\right)\tau\gamma$.
We now bound the term $\sum_{k=0}^{K-1}\|\widehat{\mathbf{d}}_{k}\|^{2}$. From (35), we have
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k+1}\|^{2}]$
$\leq(\delta+\frac{q_{1}}{1-\delta})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+\frac{q_{2}}{1-\delta}\mathbb{E}[\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}]$ (40)
$+\frac{q_{3}}{1-\delta}\mathbb{E}[\left\|\nabla F(\bar{x}_{k})\right\|^{2}]+q_{4}\sigma^{2}\leq\bar{\delta}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+q_{4}\sigma^{2}+K\mathcal{D}_{k}$ (41)
where
$\begin{split}R&\coloneqq\operatorname{\max}\{\frac{q_{3}}{1-\delta},\frac{q_{2}\tau^{2}}{1-\delta}\}.\end{split}$ (42)
Moreover, letting $\gamma\leq\min\{\bar{\gamma}_{1},\bar{\gamma}_{5}\}$, we have
$\bar{\delta}=\delta+\frac{q_{1}}{1-\delta}<1.$ (43)
Iterating (41) now gives
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]\leq\bar{\delta}_{k}\mathbb{E}[\|\widehat{\mathbf{d}}_{0}\|^{2}]+R\sum_{\ell=0}^{k-1}\bar{\delta}^{k-1-\ell}\mathcal{D}_{\ell}+\frac{q_{4}\sigma^{2}}{1-\bar{\delta}}$
and summing this inequality over $k=0,\ldots,K-1$, it follows that
$\sum_{k=0}^{K-1}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]\leq\frac{\|\widehat{\mathbf{d}}_{0}\|^{2}}{1-\bar{\delta}}+\frac{R}{1-\bar{\delta}}\sum_{k=0}^{K-1}\mathcal{D}_{k}+\frac{q_{4}\sigma^{2}K}{1-\bar{\delta}}.$ (44)
Substituting (44) into (39) and rearranging, we obtain
$\left(1-q_{5}\right)\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\frac{8\tilde{F}\left(\bar{x}^{0}\right)}{\gamma\tau}+q_{6}\|\hat{\mathbf{d}}_{0}\|^{2}+Kq_{7}\sigma^{2},$
where
$q_{5}\coloneqq\frac{c_{4}R}{1-\bar{\delta}},\quad q_{6}\coloneqq\frac{c_{4}}{1-\bar{\delta}}\quad q_{7}\coloneqq\frac{c_{4}q_{4}}{1-\bar{\delta}}+c_{5}.$ (45)
When $\gamma\leq\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}\tau^{2}}{8(\beta_{0}\kappa_{1}+\beta_{1})}$, we have $1-\bar{\delta}\geq\frac{|\tilde{\lambda}_{\max}|\rho\tau\gamma}{4},$ and $\frac{c_{4}}{1-\bar{\delta}}\leq\frac{4\kappa_{4}}{|\tilde{\lambda}_{\max}|\rho\tau}.$ Using this fact together with (42) and $\gamma\leq 1$, we have $q_{5}\leq\max\{\frac{128\kappa_{4}\tau^{2}N\beta_{0}\gamma}{|\tilde{\lambda}_{\max}|^{2}\rho^{2}\tau},\frac{24L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}N\gamma}{|\tilde{\lambda}_{\max}|^{2}\rho^{2}\tau^{2}}\}.$ Therefore, when $\gamma\leq\min\{1,\bar{\gamma}_{6},\bar{\gamma}_{7}\}$, we have
$q_{5}\leq\frac{1}{2},$ (46)
and it follows that
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\frac{16\tilde{F}\left(\bar{x}^{0}\right)}{\gamma\tau K}+\frac{2q_{6}}{K}\|\hat{\mathbf{d}}_{0}\|^{2}+2q_{7}\sigma^{2}.$ (47)
By collecting all step-size conditions, if the step-size $\gamma<\bar{\gamma}_{\text{sgd}}\coloneqq\min_{i=1,2,\ldots,7}\bar{\gamma}_{i}$, then (36), (38), (43) and (46) hold, the states $\{\mathbf{X}_{k}\}$ generated by LT-ADMM converge to the neighborhood of the stationary point, concluding the proof.
Appendix C Convergence analysis for LT-ADMM-VR
C-A Key bounds
We start by deriving an upper bound for the variance of the gradient estimator $\mathbb{E}\left[\|g_{i}\left(\phi_{i,k}^{t}\right)-\nabla f_{i}\left({\phi}_{i,k}^{t}\right)\|^{2}\right]$. Define $t_{i}^{k}$ as the averaged consensus gap of the auxiliary variables of $\{r_{i,h,k}^{k}\}_{h=1}^{m_{i}}$ at node $i$:
$t_{i,k}^{t}=\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\|r_{i,h,k}^{t}-\bar{x}_{k}\|^{2},$
$t^{t}_{k}=\sum_{i=1}^{N}t_{i,k}^{t}=\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\|\mathbf{r}_{h,k}^{t}-\bar{\mathbf{X}}_{k}\|^{2},$
${t}_{k}=\sum_{t=0}^{\tau-1}t^{t}_{k}=\sum_{t=0}^{\tau-1}\sum_{i=1}^{N}t_{i,k}^{t}.$
By the updates of $g_{i}(\phi_{i,k}^{t})$ in LT-ADMM-VR,
$\mathbb{E}\left[\|g_{i}(\phi_{i,k}^{t})-\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}\right]$
$=\mathbb{E}[\|\frac{1}{|\mathcal{B}_{i}|}\sum_{h\in\mathcal{B}_{i}}\nabla f_{i,h}\left(\phi_{i,k}^{t}\right)-\frac{1}{|\mathcal{B}_{i}|}\sum_{h\in\mathcal{B}_{i}}\nabla f_{i,h}\left(r_{i,h,k}^{t}\right)$
$-(\nabla f_{i}\left(\phi_{i,k}^{t}\right)-\frac{1}{m_{i}}\sum_{h=1}^{m_{j}}\nabla f_{i,h}(r_{i,h,k}^{k}))\|^{2}]$
$\leq\mathbb{E}\left[\|\frac{1}{|\mathcal{B}_{i}|}\sum_{h\in\mathcal{B}_{i}}\nabla f_{i,h}\left(\phi_{i,k}^{t}\right)-\frac{1}{|\mathcal{B}_{i}|}\sum_{h\in\mathcal{B}_{i}}\nabla f_{i,h}\left(r_{i,h,k}^{t}\right)\|^{2}\right]$
$\leq\mathbb{E}\left[\frac{1}{|\mathcal{B}_{i}|}\sum_{h\in\mathcal{B}_{i}}\left\|\nabla f_{i,h}\left(\phi_{i,k}^{t}\right)-\nabla f_{i,h}\left(r_{i,h,k}^{t}\right)\right\|^{2}\right]$
$=\frac{1}{|\mathcal{B}_{i}|}\sum_{h\in\mathcal{B}_{i}}\mathbb{E}\left[\left\|\nabla f_{i,h}\left(\phi_{i,k}^{t}\right)-\nabla f_{i,h}\left(r_{i,h,k}^{t}\right)\right\|^{2}\right]$
$\leq 2L^{2}\left\|\phi_{i,k}^{t}-\overline{{x}}_{k}\right\|^{2}+2L^{2}\mathbb{E}[t_{i,k}^{t}],$
where in the first inequality we use $\mathbb{E}[\|a-\mathbb{E}[a]\|^{2}]\leq\mathbb{E}[\|a\|^{2}]$ with $a=\nabla f_{i,h}(\phi_{i,k}^{t})-\nabla f_{i,h}(r_{i,h,k}^{t})$; and in the second inequality we use the smoothness of the costs. As a consequence, we have
$\mathbb{E}[\sum_{i}\sum_{t}\|g_{i}\left(\phi_{i,k}^{t}\right)-\nabla f_{i}(\phi_{i,k}^{t})\|^{2}]\leq 2L^{2}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+2L^{2}\mathbb{E}[{t}_{k}].$ (48)
Lemma 4.
Let Assumptions 1, 2 and 4 hold; when $\gamma\tau\leq\frac{2}{|\tilde{\lambda}_{\max}|\rho}$ and $6\gamma^{2}\tau(3L^{2}+\rho^{2}d_{u}^{2})<\frac{1/4}{\tau-1}$, we have
$\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]\leq\left(\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2}\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$ (49)
$+16\tau^{3}\gamma^{2}N\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]+48\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t_{k}]$
Proof.
Suppose that $\tau\geq 2$, using (31) and (48) we have
$\mathbb{E}[\left\|\Phi_{k}^{t+1}-\bar{\mathbf{X}}_{k}\right\|^{2}]$
$=\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}+\gamma\mathbf{Y}_{k}-\gamma(G(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+$
$+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}]$
$\leq\left(1+\frac{1}{\tau-1}\right)\mathbb{E}[\left\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}]+$
$+\tau\gamma^{2}\mathbb{E}[\|Y_{k}-(G(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}]$
$\leq\left(1+\frac{1}{\tau-1}\right)\mathbb{E}[\left\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}]+$
$+2\tau\gamma^{2}\mathbb{E}[\|G(\Phi_{k}^{t})-\nabla F(\Phi_{k}^{t})+\nabla F(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+$
$+\rho\mathbf{D}\Phi_{k}^{t}-\rho D\bar{\mathbf{X}}_{k}\|^{2}]+$
$+2\tau\gamma^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}+\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k}\right\|^{2}]$
$\leq\left(1+\frac{1}{\tau-1}+6\gamma^{2}\tau(2L^{2}+L^{2}+\rho^{2}d_{u}^{2})\right)\mathbb{E}[\left\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}]+$
$+2\tau\gamma^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}+\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k}\right\|^{2}]+6\tau\gamma^{2}(2L^{2}\mathbb{E}[t^{t}_{k}])$
$\leq\left(1+\frac{5/4}{\tau-1}\right)\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}]+12\tau\gamma^{2}L^{2}\mathbb{E}[t^{t}_{k}]+$
$+2\tau\gamma^{2}\mathbb{E}[\|\mathbf{Y}_{k}+\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k}\|^{2}],$ (50)
where the last inequality holds when
$6\gamma^{2}\tau(3L^{2}+\rho^{2}d_{u}^{2})\leq\frac{1/4}{\tau-1}.$ (51)
Iterating (50) for $t=0,...,\tau-1$ yields
$\mathbb{E}[\left\|\Phi_{k}^{t+1}-\bar{\mathbf{X}}_{k}\right\|^{2}]\leq\left(1+\frac{5/4}{\tau-1}\right)^{t}\mathbb{E}[\left\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\right\|^{2}]+$
$+2\tau\gamma^{2}\sum_{l=0}^{t}\left(1+\frac{5/4}{\tau-1}\right)^{l}\mathbb{E}[\|\mathbf{Y}_{k}+\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k}\|^{2}]+$
$+\sum_{l=0}^{t}\left(1+\frac{5/4}{\tau-1}\right)^{t-l}12\tau\gamma^{2}L^{2}\mathbb{E}[t_{k}^{l}]$
$\leq 4\mathbb{E}[\left\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\right\|^{2}]+8\tau^{2}\gamma^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}]$
$+48\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t^{t}_{k}],$ (52)
where in the last inequality we used $(1+\frac{a}{\tau-1})^{t}\leq\exp(\frac{at}{\tau-1})\leq\exp(a)$ for $t\leq\tau-1$. It is easy to verify that (52) also holds for $\tau=1$.
Summing (52) over $t$, and using (34), it follows that
$\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]\leq 4\tau\mathbb{E}[\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}]+8\tau^{3}\gamma^{2}\mathbb{E}[\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}$
$+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}]+48\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t_{k}]$
$\leq 4\tau\mathbb{E}[\left\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\right\|^{2}]+16\tau^{3}\gamma^{2}\mathbb{E}[\left\|\bar{\mathbf{Y}}_{k}\right\|^{2}]$
$+16\tau^{3}\gamma^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+(\rho D\bar{\mathbf{X}}_{k}-\rho D{\mathbf{X}}_{k})\right\|^{2}]$
$+48\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t_{k}]$
$\leq 4\tau\mathbb{E}[\left\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\right\|^{2}]+16\tau^{3}\gamma^{2}N\mathbb{E}[\left\|\nabla F(\bar{x}_{k})\right\|^{2}]$
$+16\tau^{3}\gamma^{2}\mathbb{E}[\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+(\rho D\bar{\mathbf{X}}_{k}-\rho D{\mathbf{X}}_{k})\|^{2}]$
$+48\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t_{k}]$
$\leq\left(\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2}\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+16\tau^{3}\gamma^{2}N\mathbb{E}[\left\|\nabla F(\bar{x}_{k})\right\|^{2}]+48\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t_{k}],$
which concludes the proof. ∎
The following lemma provides the bound on $t_{k}$.
Lemma 5.
Let $\left\{t_{k}\right\}$ be the iterates generated by LT-ADMM-VR. If $\gamma\leq\min\{\bar{\gamma}_{8},\bar{\gamma}_{9},\bar{\gamma}_{10}\}$, we have for all $k\in\mathbb{N}$:
$\mathbb{E}[t_{k}]$ $\leq 2(s_{0}+s_{1})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+2s_{2}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}],$ (53)
where
$s_{0}\coloneqq\frac{36m_{u}\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}+\frac{4\tau}{m_{l}}\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)(\gamma^{2}+\gamma)2m_{u}$
$s_{1}\coloneqq\left(\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2}\right)\frac{4m_{u}}{m_{l}}$ (54)
$s_{2}\coloneqq\frac{64m_{u}}{m_{l}}\tau^{3}\gamma^{2}N+\frac{8m_{u}}{m_{l}}(\gamma^{2}+\gamma)\tau N.$
Proof.
From Algorithm 1, $\forall k,r_{i,h,k}^{t+1}=r_{i,h,k}^{k}$ with probability $1-\frac{1}{m_{i}}$ and $r_{i,h,k}^{t+1}=\phi_{i,k}^{t+1}$ with probability $\frac{1}{m_{i}}$. Therefore, we have
$\mathbb{E}\left[t^{t+1}_{k}\right]=\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\mathbb{E}\left[\left\|\mathbf{r}_{h,k}^{t+1}-\bar{\mathbf{X}}_{k}\right\|^{2}\right]$
$=\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\mathbb{E}[(1-\frac{1}{m_{i}})\|\mathbf{r}_{h,k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}+\frac{1}{m_{i}}\|\Phi_{k}^{t+1}-\bar{\mathbf{X}}_{k}\|^{2}]$
$=\left(1-\frac{1}{m_{i}}\right)\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\mathbb{E}\left[\|\mathbf{r}_{h,k}^{t}-\bar{X}_{k}\|^{2}\right]$
$+\frac{1}{m_{i}}\mathbb{E}\left[\left\|\Phi_{k}^{t+1}-\bar{\mathbf{X}}_{k}\right\|^{2}\right].$
Now, denoting $q^{t}_{k}=\mathbf{Y}_{k}-(G(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k}),$ we have
$\left\|\Phi_{k}^{t+1}-\bar{X}_{k}\right\|^{2}=\left\|\Phi_{k}^{t+1}-\Phi_{k}^{t}+\Phi_{k}^{t}-\bar{X}_{k}\right\|^{2}$
$=\|\Phi_{k}^{t}-\bar{X}_{k}\|^{2}+2\gamma\langle\Phi_{k}^{t}-\bar{X}_{k},q^{t}_{k}\rangle+\gamma^{2}\|q^{t}_{k}\|^{2}$
$\leq(1+\gamma)\|\Phi_{k}^{t}-\bar{X}_{k}\|^{2}+(\gamma^{2}+\gamma)\|q^{t}_{k}\|^{2},$
and using (48) yields
$\mathbb{E}\left[\left\|q^{t}_{k}\right\|^{2}\right]$
$=\mathbb{E}[\|\mathbf{Y}_{k}-(\nabla F(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}]$
$+\mathbb{E}\left[\|G(\Phi_{k}^{t})-\nabla F(\Phi_{k}^{t})\|^{2}\right]$
$\leq\mathbb{E}\left[\|G(\Phi_{k}^{t})-\nabla F(\Phi_{k}^{t})\|^{2}\right]$
$+2\mathbb{E}[\|\nabla F(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\bar{\mathbf{X}}_{k}\|^{2}]$
$+2\mathbb{E}[\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}]$
$\leq\mathbb{E}\left[\|G(\Phi_{k}^{t})-\nabla F(\Phi_{k}^{t})\|^{2}\right]+2(L^{2}+\rho^{2}d_{u}^{2})\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}$
$+4\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)\|\widehat{\mathbf{d}}_{k}\|^{2}+4N\|\nabla F(\bar{x}_{k})\|^{2}$
$\leq(4L^{2}+2\rho^{2}d_{u}^{2})\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}+4\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)\|\widehat{\mathbf{d}}_{k}\|^{2}$
$+2L^{2}\mathbb{E}[t^{t}_{k}]+4N\|\nabla F(\bar{x}_{k})\|^{2}.$
Therefore we have
$\mathbb{E}\left[t^{t+1}_{k}\mid\mathcal{F}^{k}\right]=\left(1-\frac{1}{m_{i}}\right)\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\|\mathbf{r}_{h,k}^{t}-\bar{X}_{k}\|^{2}$
$+\frac{1}{m_{i}}\|\Phi_{k}^{t+1}-\bar{X}_{k}\|^{2}$
$\leq(1-\frac{1}{m_{i}})t^{t}_{k}+\frac{1}{m_{i}}\left((1+\gamma)\|\Phi_{k}^{t}-\bar{X}_{k}\|^{2}+(\gamma^{2}+\gamma)\|q^{t}_{k}\|^{2}\right)$
$\leq\left(1-\frac{1}{m_{u}}+\frac{2L^{2}}{m_{l}}(\gamma^{2}+\gamma)\right)\mathbb{E}[t^{t}_{k}]+$
$+\left(\frac{1}{m_{l}}(1+\gamma)+\frac{4L^{2}+2\rho^{2}d_{u}^{2}}{m_{l}}(\gamma^{2}+\gamma)\right)\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}+$
$+\frac{4}{m_{l}}\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)(\gamma^{2}+\gamma)\|\widehat{\mathbf{d}}_{k}\|^{2}+$
$+\frac{4N}{m_{l}}(\gamma^{2}+\gamma)\|\nabla F(\bar{x}_{k})\|^{2}$
$\leq\left(1-\frac{1}{2m_{u}}\right)\mathbb{E}[t^{t}_{k}]+\frac{2}{m_{l}}\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}+$
$+\frac{4}{m_{l}}\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)(\gamma^{2}+\gamma)\|\widehat{\mathbf{d}}_{k}\|^{2}+$
$+\frac{4N}{m_{l}}(\gamma^{2}+\gamma)\|\nabla F(\bar{x}_{k})\|^{2}$ (55)
where the last inequality holds when
$\frac{2L^{2}}{m_{l}}(\gamma^{2}+\gamma)<\frac{1}{2m_{u}},\quad\gamma+(4L^{2}+2\rho^{2}d_{u}^{2})(\gamma^{2}+\gamma)<1,$ (56)
which can be satisfied by choosing $\gamma<\min\{\bar{\gamma}_{8},\bar{\gamma}_{9}\}$. Iterating (55) for $t=0,...,\tau-1$ then yields:
$\mathbb{E}[t^{t}_{k}]\leq(1-\frac{1}{2m_{u}})^{t}\mathbb{E}[\left\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\right\|^{2}]$
$+\frac{4}{m_{l}}\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)(\gamma^{2}+\gamma)\sum_{l=0}^{t-1}(1-\frac{1}{2m_{u}})^{l}\|\widehat{\mathbf{d}}_{k}\|^{2}$
$+\frac{4N}{m_{l}}(\gamma^{2}+\gamma)\sum_{l=0}^{t-1}(1-\frac{1}{2m_{u}})^{l}\|\nabla F(\bar{x}_{k})\|^{2}$
$+\frac{2}{m_{l}}\sum_{l=0}^{t-1}(1-\frac{1}{2m_{u}})^{t-1-l}\mathbb{E}[\|\Phi_{k}^{l}-\bar{\mathbf{X}}_{k}\|^{2}$
$\leq 9\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}(1-\frac{1}{2m_{u}})^{t}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+\frac{4N}{m_{l}}(\gamma^{2}+\gamma)\sum_{l=0}^{t-1}(1-\frac{1}{2m_{u}})^{l}\|\nabla F(\bar{x}_{k})\|^{2}$
$+\frac{2}{m_{l}}\sum_{l=0}^{t-1}(1-\frac{1}{2m_{u}})^{t-1-l}\mathbb{E}[\|\Phi_{k}^{l}-\bar{\mathbf{X}}_{k}\|^{2}$
$+\frac{4}{m_{l}}\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)(\gamma^{2}+\gamma)2m_{u}\|\widehat{\mathbf{d}}_{k}\|^{2}.$
Summing the above relation over $t=0,1,...,\tau-1$ we get:
$\mathbb{E}[t_{k}]$
$\leq\left(\frac{36m_{u}\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}+\frac{4\tau}{m_{l}}(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})(\gamma^{2}+\gamma)2m_{u}\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+\frac{4m_{u}}{m_{l}}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\frac{4N}{m_{l}}(\gamma^{2}+\gamma)2m_{u}\tau\|\nabla F(\bar{x}_{k})\|^{2},$
and using (49) then yields
$\mathbb{E}[t_{k}]$ $\leq(s_{0}+s_{1})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+s_{2}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]+$
$+\frac{192m_{u}}{m_{l}}\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t_{k}],$
where $s_{0}$, $s_{1}$ and $s_{2}$ are defined in (54).
Letting $\gamma<\bar{\gamma}_{10}$, it follows that
$\frac{192m_{u}}{m_{l}}\tau^{2}\gamma^{2}L^{2}<\frac{1}{2},$ (57)
and thus $\mathbb{E}[t_{k}]\leq 2(s_{0}+s_{1})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+2s_{2}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}].$
Lemma 6.
Let Assumptions 1, 2 and 4 hold; when $\gamma<\min\{\bar{\gamma}_{1},\bar{\gamma}_{11}\}$, it holds that $\forall k\geq 0$
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k+1}\|^{2}]\leq(\delta+\frac{\tilde{q}_{1}}{1-\delta})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$ (58)
$+\frac{\tilde{q}_{2}}{1-\delta}\mathbb{E}[\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}]+\frac{\tilde{q}_{3}}{1-\delta}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]+\tilde{q}_{4}\mathbb{E}[t_{k}],$
where
$\delta\coloneqq 1-\frac{|\tilde{\lambda}_{\max}|\rho\tau\gamma}{2},\tilde{c}_{0}\coloneqq\gamma^{2}\tau\tilde{\beta}_{0},\tilde{c}_{1}\coloneqq\gamma^{3}\tau^{2}\tilde{\beta}_{1},$
$\tilde{c}_{2}\coloneqq{12L^{2}\gamma^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}N,\tilde{c}_{3}\coloneqq\gamma^{2}\tau\tilde{\beta}_{3},$
$\tilde{q}_{1}:=\tilde{c}_{0}\gamma\tau\kappa_{1}+\tilde{c}_{1},\tilde{q}_{2}\coloneqq\tilde{c}_{2},\quad\tilde{q}_{3}:=16\tau^{3}\gamma^{2}N\tilde{c}_{0},$
$\tilde{q}_{4}:=\left(48\tau^{2}\gamma^{2}L^{2}\tilde{c}_{0}+\tilde{c}_{3}\right)/(1-\delta).$
Proof.
When $\gamma<\min\{\bar{\gamma}_{1},\bar{\gamma}_{11}\}$ we have
$\gamma\tau\leq\frac{2}{|\tilde{\lambda}_{\max}|\rho},\quad|\tilde{\lambda}_{\min}|\rho\tau\gamma<2,$ (59)
$6\gamma^{2}\tau(3L^{2}+\rho^{2}d_{u}^{2})<\frac{1/4}{\tau-1}.$
Substituting (48) into (28) then yields
$\|\widehat{\mathbf{h}}_{k}\|^{2}\leq\tilde{c}_{0}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\tilde{c}_{1}\|\widehat{\mathbf{d}}_{k}\|^{2}+\tilde{c}_{2}\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}+\tilde{c}_{3}t_{k},$
and together with (49) and (17), it proves that (58) holds. ∎
C-B Theorem 2
Based on (30), substituting $y=\bar{x}_{k+1}$ and $z=\bar{x}_{k}$ into (37) and using (27), (48), we get
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]$
$\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\gamma\mathbb{E}[\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}(\sum_{t}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)+\epsilon_{k})\rangle]$
$+2\gamma^{2}\tau L\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\frac{2\gamma^{2}\tau L}{N}(2L^{2}\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]+2L^{2}\mathbb{E}[{t}_{k}])$
$+\frac{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau(\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]+\tau\mathbb{E}[\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}])}{N}.$
Similarly to Appendix B-B, using $2\langle a,b\rangle=\|a\|^{2}+\|b\|^{2}-\|a-b\|^{2}$ we have
-- $\mathbb{E}[\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}\sum_{t}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)+\frac{\epsilon_{k}}{N}\rangle]$
$\leq-\frac{\tau}{4}\mathbb{E}[\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}]-\frac{1}{2}\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\left(\frac{L^{2}}{2N}+\frac{2\rho^{2}d_{u}^{2}}{N}\right)\mathbb{E}[\widehat{\mathbf{\Phi}}_{k}\|^{2}]+\frac{2\rho^{2}d_{u}^{2}\tau}{N}\mathbb{E}[\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}].$
Combining the two inequalities above, and using (16), yields
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\frac{\gamma\tau}{4}\mathbb{E}[\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}]$ (60)
$-\frac{\gamma}{2}(1-4\gamma\tau L)\mathbb{E}\left[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}\right]$
$+\left(\frac{\gamma L^{2}}{2N}+\frac{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}+4\gamma^{2}\tau L^{3}}{N}\right)\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]$
$+\frac{18\gamma\rho^{2}d_{u}^{2}\tau+18\gamma^{2}\rho^{2}d_{u}^{2}\tau^{2}L}{N}\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+\frac{4\gamma^{2}\tau L^{3}\mathbb{E}[t_{k}]}{N}.$
Combining (60) together with (49) and (53), and using $\gamma\leq 1$, yields
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\frac{\gamma\tau}{4}\mathbb{E}[\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}]$
$-\frac{\gamma}{2}(1-4\gamma\tau L)\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\left(k_{0}k_{2}+2(k_{3}k_{0}+k_{4})s_{2}\right)\mathbb{E}[\|\nabla F(x)\|^{2}]$
$+\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}],$
where
$k_{0}\coloneqq\alpha_{0}\gamma,k_{1}\coloneqq\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2}$
$k_{2}\coloneqq 16\tau^{3}\gamma^{2}N,k_{3}\coloneqq 48\tau^{2}\gamma^{2}L^{2}$
$k_{4}\coloneqq\frac{4\gamma^{2}\tau L^{3}}{N},k_{5}\coloneqq\frac{36\gamma^{2}\rho d_{u}^{2}\tau^{2}+36\gamma^{3}\rho d_{u}^{2}\tau^{3}L}{N|\tilde{\lambda}_{\max}|}.$
Letting $\gamma\leq\min\{1,\bar{\gamma}_{12},\bar{\gamma}_{13}\}$, then
$k_{0}k_{2}+2(k_{3}k_{0}+k_{4})s_{2}\leq\frac{\gamma\tau}{8},\quad 4\gamma\tau L\leq\frac{3}{4},$ (61)
and we can upper bound the previous inequality by
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\frac{\gamma\tau}{8}\mathbb{E}[\|\nabla F\left(\bar{x}_{k}\right)\|^{2}]$
$-\frac{\gamma}{8}\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}].$
Rearranging we get
$\mathcal{D}_{k}$ $\leq\frac{8}{\gamma\tau}(\mathbb{E}[\tilde{F}\left(\bar{x}_{k}\right)]-\mathbb{E}[\tilde{F}\left(\bar{x}_{k+1}\right)])$ (62)
$+\frac{8}{\gamma\tau}\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}],$
where $\mathcal{D}_{k}$ is defined in (7), and $\tilde{F}\left(\bar{x}_{k}\right)=F\left(\bar{x}_{k}\right)-F(x^{*})$. Summing (62) over $k=0,1,\ldots,K-1$ and using $-\tilde{F}\left(\bar{x}_{k}\right)\leq 0$, it holds that
$\sum_{r=0}^{K-1}\mathcal{D}_{k}\leq\frac{8\tilde{F}\left(\bar{x}_{0}\right)}{\gamma\tau}$ (63)
$+\frac{8\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)}{\gamma\tau}\sum_{k=0}^{K-1}\|\mathbb{E}[\hat{\mathbf{d}}_{k}\|^{2}].$
According to (58) and (53), we derive that $\forall k\geq 0$,
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k+1}\|^{2}]\leq\left(\delta+\frac{\tilde{q}_{1}}{1-\delta}+2\tilde{q}_{4}(s_{0}+s_{1})\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$ (64)
$+\frac{\tilde{q}_{2}}{1-\delta}\mathbb{E}[\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}]+\left(\frac{\tilde{q}_{3}}{1-\delta}+2\tilde{q}_{4}s_{2}\right)\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]$
$\leq\tilde{\delta}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+\tilde{R}\mathcal{D}_{k},$
where
$\tilde{R}=\max\left\{\frac{\tilde{q}_{2}\tau^{2}}{1-\delta},\frac{\tilde{q}_{3}}{1-\delta}+2\tilde{q}_{4}s_{2}\right\}$
$=\max\left\{\frac{2\tilde{q}_{2}\tau}{\rho\gamma|\tilde{\lambda}_{max}|},\frac{2(\tilde{q}_{3}+2(48\tau^{2}\gamma^{2}L^{2}\tilde{c}_{0}+\tilde{c}_{3})s_{2})}{\tau\rho\gamma|\tilde{\lambda}_{max}|}\right\}.$
Letting $\gamma\leq\min\{\bar{\gamma}_{1},\bar{\gamma}_{14}\}$, then
$\tilde{\delta}=\delta+\frac{\tilde{q}_{1}}{1-\delta}+2\tilde{q}_{4}(s_{0}+s_{1})<1.$ (65)
Iterating (64) yields $\forall k\geq 1$, $\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]\leq\tilde{\delta}^{k}\mathbb{E}[\|\widehat{\mathbf{d}}_{0}\|^{2}]+\tilde{R}\sum_{\ell=0}^{k-1}\tilde{\delta}^{k-1-\ell}\mathcal{D}_{\ell},$ and summing over $k=0,\ldots,K-1$ it holds that
$\sum_{k=0}^{K-1}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]\leq\frac{1}{1-\tilde{\delta}}\|\widehat{\mathbf{d}}_{0}\|^{2}+\sum_{k=0}^{K-1}\frac{\tilde{R}}{1-\tilde{\delta}}\mathcal{D}_{k}.$ (66)
Substituting (66) into (63), and rearranging, we obtain
$\left(1-\frac{\tilde{R}}{1-\tilde{\delta}}\frac{8}{\gamma\tau}\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)\right)\sum_{k=0}^{K-1}\mathcal{D}_{k}$
$\leq\frac{8\tilde{F}\left(\bar{x}_{0}\right)}{\gamma\tau}+\frac{8\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)}{\gamma\tau(1-\tilde{\delta})}\|\widehat{\mathbf{d}}_{0}\|^{2}.$
Similarly to the previous analysis, when $\gamma\leq\frac{\gamma_{14}}{2}=\bar{\gamma}_{15}$, $1-\tilde{\delta}\geq\frac{|\tilde{\lambda}_{\max}|\rho\tau\gamma}{4}.$ Let $\gamma\leq\min\{\bar{\gamma}_{16},\bar{\gamma}_{17}\}$, then
$\frac{\tilde{R}}{1-\tilde{\delta}}\frac{8}{\gamma\tau}\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)\leq\frac{1}{2},$
and therefore it follows that
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\frac{16\tilde{F}\left(\bar{x}_{0}\right)}{K\gamma\tau}$
$+\frac{16}{K\gamma\tau}\left(k_{5}+k_{0}k_{1}+2(k_{3}k_{0}+k_{4})(s_{0}+s_{1})\right)\frac{1}{1-\tilde{\delta}}\|\widehat{\mathbf{d}}_{0}\|^{2}.$
By collecting all step-size conditions, if the step-size $\gamma$ satisfies (51), (56), (57), (59), (61), (65) and (C-B), then the states $\{\mathbf{X}_{k}\}$ generated by LT-ADMM-VR converge to the stationary point, concluding the proof.